### v1

# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

Abstract
In was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution. However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems. In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times. Moreover, our methods are considerably easier to implement.
1 Introduction
Convex optimization techniques have been extremely successful in problems where the goal is to recover a sparse signal given certain measurements, e.g., [4, 5, 6, 7, 8, 9, 17, 18]. This work concerns certain canonical and popular convex relaxations to continuous optimization problems in which the goal is to find some optimal solution which is assumed to be sparse. We begin by presenting our two main examples.
Example 1: $\ell_{1}$-constrained minimization for entry-wise sparse recovery.
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~{}\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (1)
where $f$ is convex and continuously differentiable, and we assume that Problem (1) indeed admits a unique optimal solution ${\mathbf{x}}^{*}$ such that $\textrm{nnz}({\mathbf{x}}^{*})<<n$, where $\textrm{nnz}(\cdot)$ denotes the number of non-zero entries.
Example 2: trace-constrained minimization for low-rank matrix recovery.
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{}{\mathbf{X}}\succeq 0,~{}\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}),$ (2)
where $\mathbb{S}^{n}$ is the space of real symmetric $n\times{}n$ matrices and the relation ${\mathbf{X}}\succeq 0$ denotes ${\mathbf{X}}$ is positive semidefinite. Here we also assume $f$ is convex and continuously differentiable, and that Problem (2) indeed admits a unique optimal solution ${\mathbf{X}}^{*}$ such that $\textrm{rank}({\mathbf{X}}^{*})<<n$.
[…]
In recent works [16, 14], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that the optimal solution is indeed sparse/low-rank, it is possible to obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipchitz continuity constant of the gradient w.r.t. the $\ell_{1}$ norm, $\alpha_{2}$ is the quadratic growth constant w.r.t. the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of the optimal solution. This condition number could indeed be far better than the pure $\ell_{2}$ or pure $\ell_{1}$ condition numbers. For instance, in the context of Problem (1), consider the case that $f$ is convex quadratic of the following very simple form:
$f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}{\mathbf{A}}({\mathbf{x}}-{\mathbf{x}}^{*});\quad{\mathbf{A}}:=\alpha{\mathbf{I}}+\mathbf{1}\mathbf{1}^{\top},$ (3)
where $\mathbf{1}$ denotes the all-ones vector in $\mathbb{R}^{n}$ and $\alpha$ is a positive scalar.
Indeed if $\|{{\mathbf{x}}^{*}}\|_{1}\leq R$, ${\mathbf{x}}^{*}$ is the unique optimal solution to Problem (1) for this $f$. One can easily observe that the Euclidean condition number for this problem is given by $\frac{\beta_{2}}{\alpha_{2}}=\frac{\lambda_{\max}({\mathbf{A}})}{\lambda_{\min}({\mathbf{A}})}=\frac{n+\alpha}{\alpha}$. On the other-hand, a simple calculation shows that the $\ell_{1}$-Lipchitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$. To bound the $\ell_{1}$-quadratic growth constant we observe that due the the structure of the matrix ${\mathbf{A}}$,
$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$ (4)
In particular, assuming ${\mathbf{x}}^{*}\geq 0$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., ${\mathbf{x}}^{*}$ lies on the boundary of the $\ell_{1}$ ball, as often expected) we have that
$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~{}|~{}{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=R\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
Thus, in worst-case, the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
However, in the example above, if ${\mathbf{x}}^{*}$ satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
The example in Eq. (4) could be directly extended also to Problem (9) by taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\textrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$, where $\|{\cdot}\|_{2}$ here is the Euclidean norm in $\mathbb{S}^{n}$, and instead of the entry-wise $\ell_{1}$ norm we consider the Schatten 1-norm (i.e., the $\ell_{1}$ norm w.r.t. the eigenvalues), and $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$.
The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t. unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization modelindeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t. the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsity intermediate iterates, see for instance Corollary 5 in [14].
While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is large. For instance, in the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next. In the case of Problem (2) the time required by these methods to project onto the positive semidefinite cone ($O(n^{3})$ in standard implementations) may be prohibitive.
In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the improved condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates which depend only on its square root). The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solution), which often allows for significant improvements in the runtime of each iteration. In particular, our algorithm updates its iterates similarly to the classical Frank-Wolfe algorithm [15]:
${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}\quad\textrm{for some}\quad\eta\in[0,1],$ (5)
where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for ${\mathbf{x}}^{*}$.
For instance, in case of Problem (1) with a quadratic objective function $f({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the fact that ${\mathbf{v}}_{t}$ is $s$-sparse, reduces the time to update the gradient direction from $O(n^{2})$ to $O(sn)$. In the case of Problem (2), since $\textrm{rank}({\mathbf{v}}_{t})\leq s$, computing it won’t require a full eigen-decomposition ($O(n^{3})$ time in standard implementations), but only computing the $s$ leading components in the eigen-decomposition, which is far more efficient.
Two additional advantages which make our approach considerably simpler (compared to [16, 14]) are: I. it does not require restarts, i.e., it is a single-loop algorithm, and II. the updates in our algorithm (the way ${\mathbf{v}}_{t}$ in Eq. (5) is computed) are Euclidean, which often leads to significantly simpler computations.
Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively. We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant). Additionally, in case of the matrix Problem (2), we shall have an additional term in the condition number, namely the condition number will be of the form: $\left({s+\frac{R^{2}}{\lambda_{\min}({\mathbf{X}}^{*})^{2}}}\right)\frac{\beta_{1}}{\alpha_{2}}$, where $\lambda_{\min}({\mathbf{X}}^{*})$ denotes the smallest non-zero eigenvalue of the optimal solution.
The rest of the paper is organized as follows. In Section 2 we present a unified setup and optimization problem which captures the spare convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis. In Section 3 we instantiate our generic algorithm in the special cases of Problems (1) and (2). In Section 4 we give a simple numerical demonstration for the performance of our method on Problem (1). Finally, in Section 5 we outline several open questions.
2 Unified Setup, Algorithm, and Convergence Analysis
Our formal unified setup for tackling the sparse convex optimization problems under consideration involves the following convex optimization problem:
$\min_{{\mathbf{x}}\in\mathcal{K}}f({\mathbf{x}}),$ (6)
[…]
$\forall{\mathbf{x}},{\mathbf{y}}\in\mathcal{K}:\quad\|{\nabla{}f({\mathbf{x}})-\nabla{}f({\mathbf{y}})}\|_{*}\leq\beta\|{{\mathbf{x}}-{\mathbf{y}}}\|,$
where $\|{\cdot}\|_{*}$ denotes the dual norm.
We let $\|{{\mathbf{x}}}\|_{2}=\sqrt{\langle{{\mathbf{x}},{\mathbf{x}}}\rangle}$ denote the Euclidean norm.
We shall assume Problem (6) admits a unique optimal solution ${\mathbf{x}}^{*}$, and satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
$\forall{\mathbf{x}}\in\mathcal{K}:\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (7)
We now make an additional assumption that further specializes Problem (6). In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for the optimal solution ${\mathbf{x}}^{*}$.
Assumption 1.
There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such ${\mathbf{x}}^{*}\in\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
  1. 1.
    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
  2. 2.
    for all ${\mathbf{x}}\in\mathcal{K}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
  3. 3.
    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathcal{K}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors which are $s$-sparse (in some sense), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to $s$-sparse vectors. Additionally, while not part of Assumption 1, we shall implicitly assume that it is computationally efficient to compute a Euclidean projection onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$ (in case the projection is not a singleton, compute one element in the set of possible projections). Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively.
We can now present our algorithm for solving the generic Problem (6) (under the assumptions listed above). Our algorithm is composed of three simple ingredients. First, we use a Frank-Wolfe like scheme [15] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq. (5). Second, and differently from the classical Frank-Wolfe method in which the new point ${\mathbf{v}}_{t}$ is the output of a linear optimization oracle, here we take it to be a Euclidean proximal gradient step w.r.t. to restricted set $\mathcal{K}_{s}$ (instead of w.r.t. the original set $\mathcal{K}$). This is similar to approach taken in previous works [1, 10, 12, 13] (also referred to as a weak proximal oracle [10, 13]). Third, the $\ell_{2}$-term in this proximal gradient step is not w.r.t. to the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
Algorithm 1 Frank-Wolfe with sparse proximal oracle and hard-thresholding
  ${\mathbf{x}}_{1}\leftarrow$ some arbitrary point in $\mathcal{K}$
  for $t=1,2\dots$ do
     $\widehat{{\mathbf{x}}}_{t}\leftarrow\Pi_{\mathbb{E}_{s}}[{\mathbf{x}}_{t}]$
     ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{2s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
     ${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}$
  end for
Theorem 1.
Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with step-size $\eta=\frac{\alpha_{2}}{4\beta(4s+s^{*}_{\mathcal{K}})}$. Then,
$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(4s+s^{*}_{\mathcal{K}})}}\right).$
Proof.
Fix some iteration $t$ of the algorithm. The Lipchitz continuity of the gradient implies the following well known inequality:
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\frac{\beta\eta^{2}}{2}\|{{\mathbf{v}}_{t}-{\mathbf{x}}_{t}}\|^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Since $\widehat{{\mathbf{x}}}_{t}$ and ${\mathbf{v}}_{t}$ are in $\mathbb{E}_{s}$, using Item 1 in Assumption 1 we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of ${\mathbf{v}}_{t}$ in the algorithm and the assumption that ${\mathbf{x}}^{*}\in\mathcal{K}_{s}$ we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm, Item 3 and then Item 2 in Assumption 1 we have that,
$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
Combining the last two inequalities we have that,
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(2s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+2\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm.
Using both the convexity and quadratic growth of $f$ (Eq. (7)), we have that
$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right)+\frac{2\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right).$
Subtracting $f({\mathbf{x}}^{*})$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
3 Applications
3.1 $\ell_{1}$-ball setup
We recall the first problem in our Introduction:
$\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~{}\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (8)
[…]
We shall require the following simple lemma (a proof is brought in the appendix for completeness).
Lemma 1.
Let ${\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{n}$ such that $\|{{\mathbf{y}}}\|_{1}\leq\|{{\mathbf{x}}}\|_{1}$. Then,
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\textrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Let $s$ be such that $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~{}|~{}\textrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~{}|~{}\|{{\mathbf{x}}}\|_{1}\leq R,~{}\textrm{nnz}({\mathbf{x}})\leq s\}$. In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=2s$.
Note that computing the hard-thresholding step in Algorithm 1 (the point $\widehat{{\mathbf{x}}}_{t}$) is straightforward by zeroing-out all but the largest (in absolute value) entries in ${\mathbf{x}}_{t}$.
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$ onto the $s$-sparsified $\ell_{1}$-ball $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the $n-s$ lower entries in ${\mathbf{z}}_{t}$) to obtain again a point the $n$-dimensional $\ell_{1}$-ball of radius $R$. See [3] for proof of the correctness of this procedure.
Overall, aside from the gradient computation, each iteration of Algorithm 1 requires $O(n\log{}s)$ time (which is the time to find the $s$ largest entries in an array of length $n$ — the most time consuming operation).
Corollary 1.
When applied to Problem (8) with an optimal solution ${\mathbf{x}}^{*}$ which satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of gradient computations, the runtime of each iteration is $O(n\log{s})$. In particular, if $f$ is (convex) quadratic, the overall runtime of each iteration is $O(sn)$.
Remark 1.
It is straightforward to extend the result above to the case in which ${\mathbf{x}}$ is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~{}|~{}\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as ${\mathbf{x}}^{*}$ satisfies $\|{{\mathbf{x}}}\|_{1}=R$. For instance when $\mathcal{C}=\{{\mathbf{x}}~{}|~{}{\mathbf{x}}\geq 0,\sum_{i=1}^{n}{\mathbf{x}}_{i}=R\}$, we have that $\mathcal{K}$ is simply a scaled simplex for which the assumption $\|{{\mathbf{x}}^{*}}\|_{1}=R$ always holds. Moreover, when $\mathcal{C}$ is closed under permutation of coordinates (e.g., an $\ell_{p}$ ball or its non-negative restriction), the sparse proximal step (computation of ${\mathbf{v}}_{t}$ in Algorithm 1) admits efficient implementation as detailed in [3].
3.2 Bounded positive semidefinite cone setup
We recall the second problem in our Introduction:
$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{}{\mathbf{X}}\succeq 0,~{}\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}).$ (9)
[…]
The following lemma is analogues to Lemma 1. A proof is given in the appendix for completeness.
Lemma 2.
Let ${\mathbf{X}},{\mathbf{Y}}\in\{{\mathbf{Z}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{Z}}\succeq 0\}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then,
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{\textrm{rank}({\mathbf{X}})}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{\min}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$
where $\lambda_{\min}({\mathbf{X}})>0$ denotes the smallest non-zero eigenvalue of ${\mathbf{X}}$.
Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{X}}\succeq 0,~{}\|{{\mathbf{X}}}\|_{1}\leq R,~{}\textrm{rank}({\mathbf{X}})\leq s\}$. In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=\left({2\sqrt{s}+\frac{2R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$ (note that for any feasible ${\mathbf{Y}}$ it holds that $\|{{\mathbf{Y}}}\|_{2}\leq R$).
Here the hard-thresholding step in Algorithm 1 (computing the point $\widehat{{\mathbf{x}}}_{t}$) is done by computing the rank-$s$ approximation of the current iterate ${\mathbf{x}}_{t}$, which requires computing only the top $s$ components in the SVD of ${\mathbf{x}}_{t}$ (and discarding the rest).
Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest signed eigenvalues) in the eigen-decomposition of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, ${\mathbf{u}}_{i}\in\mathbb{R}^{n},i=1,\dots,s$. Then, project the vector of eigenvalues $(\lambda_{1},\dots,\lambda_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~{}|~{}{\mathbf{y}}\geq 0,~{}\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of eigenvalues $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$. Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ (i.e., the eigenvectors remain unchanged). For the correctness of this procedure see Lemma 3 in the appendix.
Overall, aside from the gradient computation, the runtime of each iteration of Algorithm 1 is dominated by computing two rank-$s$ eigen-decompositions of a $n\times n$ symmetric matrix, which is far more efficient than a full-rank eigen-decomposition once $s<<n$.
Corollary 2.
When applied to Problem (9) with an optimal solution ${\mathbf{X}}^{*}$ which satisfies $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and $\textrm{Tr}({\mathbf{x}}^{*})=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\left({\sqrt{s}+\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}\frac{\beta_{1}}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of the gradient computations, the runtime of each iteration is dominated by two rank-$s$ eigen-decompositions of $n\times n$ symmetric matrices.
Comparing the rate in Corollary 2 to the state-of-the-art non-Euclidean accelerated gradient method in [14], which has convergence rate $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$, we can see that our rate is not only unaccelerated (i.e., linear in the condition number $\frac{s\beta_{1}}{\alpha_{2}}$ as opposed to linear in its square-root), but we also have an additional term of $\left({\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$, which may be significantly larger than 1. On the other-hand, the method in [14] requires a full eigen-decomposition of a $n\times n$ matrix each iteration, which is often prohibitive for large $n$.
Remark 2.
As in Remark 1, here also we can extend Problem (9) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{X}}\succeq 0,~{}\textrm{Tr}({\mathbf{X}})\leq R\}$, where $\mathcal{C}\subset\mathbb{S}^{n}$ is closed and convex, as long as $\textrm{Tr}({\mathbf{X}}^{*})=R$. In particular, when $\mathcal{C}$ is a ball w.r.t. a Schatten $p$-norm ($\ell_{p}$ norm w.r.t. the eigenvalues), one can extend Lemma 3 accordingly.
4 Some Numerical Evidence
We present simple numerical simulations to demonstrate the benefit of the improved convergence rate of our method, which depends on a mixed-norm condition number, in the context of the $\ell_{1}$-constrained Problem (1). Similarly to the example in Eq. (4), we set $f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}({\mathbf{I}}+3\cdot\mathbf{1}\mathbf{1}^{\top})({\mathbf{x}}-{\mathbf{x}}^{*})$, where $\mathbf{1}$ denotes the all-ones vector. For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipchitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively. We compare the convergence rate of our Algorithm 1 to a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]).
We set the dimension $n=3000$, the radius of the $\ell_{1}$ ball to $R=10$. The optimal solution ${\mathbf{x}}^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to ±$\pm\frac{R}{\textrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\textrm{nnz}({\mathbf{x}}^{*})$ is set in advanced, and for each level of sparsity we run 10 i.i.d. experiments (where in each one we resample ${\mathbf{x}}^{*}$). Accordingly, the figure below presents the average approximation error over 10 i.i.d. experiments.
When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$. Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$. Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in $(0,1)$). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
From Figure 1, we can clearly see that Algorithm 1 is significantly faster for highly-sparse instances, strongly outperforming even a state-of-the-art accelerated method, while as ${\mathbf{x}}^{*}$ becomes less-sparse, this advantage vanishes.
(a) $\textrm{nnz}({\mathbf{x}}^{*})=10$
(b) $\textrm{nnz}({\mathbf{x}}^{*})=30$
(c) $\textrm{nnz}({\mathbf{x}}^{*})=50$
Figure 1: Comparison of Algorithm 1 and the V-FISTA algorithm for Problem (1).
5 Open Questions
We conclude with several open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{s\beta_{1}}{\alpha_{2}}}$ while maintaining the sparse updates. Finally, when comparing the matrix Problem (9) with Problem (8), it is interesting if the additional term of $R^{2}/\lambda_{\min}({\mathbf{X}}^{*})^{2}$ in the condition number could be removed.
Acknowledgments
This work was funded by the European Union (ERC, ProFreeOpt, 101170791). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
We would like to thank Amir Beck and Marc Teboulle for fruitful discussions during the early stages of this work.
Appendix A Proof of Lemma 1
Proof.
Denote $k=\textrm{nnz}({\mathbf{x}})$. Let $S\subseteq[n]$ be the set of indices for non-zero entries in ${\mathbf{x}}$. Let ${\mathbf{x}}_{1},{\mathbf{y}}_{1}\in\mathbb{R}^{k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries indexed by $S$, respectively. Accordingly, let ${\mathbf{x}}_{2},{\mathbf{y}}_{2}\in\mathbb{R}^{n-k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries not indexed by $S$. It holds that,
$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}$ $=\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}_{2}}\|_{1}$
$\leq\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}_{2}}\|_{1}$
$=\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{y}}}\|_{1}-\|{{\mathbf{y}}_{1}}\|_{1}$
$\leq\ \|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}+\|{{\mathbf{x}}_{1}}\|_{1}-\|{{\mathbf{y}}_{1}}\|_{1}$
$\leq 2\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{1}$
$\leq 2\sqrt{k}\|{{\mathbf{x}}_{1}-{\mathbf{y}}_{1}}\|_{2}$
$\leq 2\sqrt{k}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
Appendix B Proof of Lemma 2
Proof.
Denote $k=\textrm{rank}({\mathbf{X}})$. Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$. Also, let ${\mathbf{Q}}\in\mathbb{S}^{n}$ denote the projection matrix onto subspace spanned by the top $k$ eigenvectors of ${\mathbf{Y}}$ (corresponding to the $k$ largest signed eigenvalues). Recall that $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm and Euclidean norm in $\mathbb{S}^{n}$, respectively. In the following we also let $\sigma_{i}({\mathbf{Z}})$ denote the $i$th largest singular value of some real matrix ${\mathbf{Z}}$. It holds that,
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp})({\mathbf{X}}-{\mathbf{Y}})}\|_{1}$
$=\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})+{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{P}}{\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(a)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(b)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\sqrt{k}\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\leq 2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
$\underset{(d)}{\leq}2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{2}\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2},$
where (a) follows from the assumption of the lemma, (b) follows since on one hand $\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, and on the other hand, using von Neumann’s inequality for the singular values (see for instance [19]), we have that $\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}}{\mathbf{Y}})\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}})\sigma_{i}({\mathbf{Y}})\leq\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, where we have used the fact that $\textrm{rank}({\mathbf{P}})\leq k$ and $\|{{\mathbf{P}}}\|_{\infty}=1$. Inequality (c) follows since ${\mathbf{X}}={\mathbf{P}}{\mathbf{X}}$, and (d) follows again from von Neumann’s inequality and the Cauchy-Schwarz inequality:
$\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$ $\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})\sigma_{i}({\mathbf{Y}})$
$\leq\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})^{2}}\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Y}})^{2}}$
$=\|{{\mathbf{Y}}}\|_{2}\cdot\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}.$
Using the Davis-Kahan Theorem (see for instance Lemma A.1 in [11]) we have that,
$\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}\leq\frac{2\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}.$
Thus, we conclude that
$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{k}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
Appendix C Sparse Projections onto the Bounded Positive Semidefinite Cone
Lemma 3.
Let ${\mathbf{Z}}\in\mathbb{S}^{n}$ with eigen-decomposition ${\mathbf{Z}}=\sum_{i=1}^{n}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ ($\lambda_{1}\geq\lambda_{2}\geq...$) and $s\in[n]$. A Euclidean projection of ${\mathbf{X}}$ onto $\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{X}}\succeq 0,~{}\textrm{Tr}({\mathbf{X}})\leq R,\textrm{rank}({\mathbf{X}})\leq s\}$ is given by the matrix ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, where $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of the vector $(\lambda_{1},\dots,\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~{}|~{}{\mathbf{x}}\geq 0,~{}\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
Proof.
[…]
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}$ $=\|{{\mathbf{Y}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}}\|_{2}^{2}-2\langle{{\mathbf{Y}}-{\mathbf{Z}}^{\prime},{\mathbf{Z}}}\rangle$
$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i}$
$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\lambda_{i}+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i},$
where the last inequality is due to von Neumann’s trace inequality.
Rearranging, we have that
$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\lambda_{i})^{2}-\sum_{j=1}^{s}(\lambda_{j}^{\prime}-\lambda_{j})^{2}\geq 0,$
where the last inequality follows since $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of $(\lambda_{1},\dots\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~{}|~{}{\mathbf{x}}\geq 0,~{}\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$. ∎

### v2

@@
 Abstract
-In was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution.
+It was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution.
 However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems.
@@
 $\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (1)
-where $f$ is convex and continuously differentiable, and we assume that Problem (1) indeed admits a unique optimal solution ${\mathbf{x}}^{*}$ such that $\textrm{nnz}({\mathbf{x}}^{*})<<n$, where $\textrm{nnz}(\cdot)$ denotes the number of non-zero entries.
+[…]
 Example 2: trace-constrained minimization for low-rank matrix recovery.
@@
 To bound the $\ell_{1}$-quadratic growth constant we observe that due the the structure of the matrix ${\mathbf{A}}$,
-$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$ (4)
+$\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$
 In particular, assuming ${\mathbf{x}}^{*}\geq 0$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., ${\mathbf{x}}^{*}$ lies on the boundary of the $\ell_{1}$ ball, as often expected) we have that
@@
 unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization model indeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t.
-the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsity intermediate iterates, see for instance Corollary 5 in [14].
+the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsify intermediate iterates, see for instance Corollary 5 in [14].
 While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is large.
@@
 We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant).
-Additionally, in case of the matrix Problem (2), we shall have an additional term in the condition number, namely the condition number will be of the form: $\left({s+\frac{R^{2}}{\lambda_{\min}({\mathbf{X}}^{*})^{2}}}\right)\frac{\beta_{1}}{\alpha_{2}}$, where $\lambda_{\min}({\mathbf{X}}^{*})$ denotes the smallest non-zero eigenvalue of the optimal solution.
 The rest of the paper is organized as follows.
@@
 Let ${\mathbf{X}},{\mathbf{Y}}\in\{{\mathbf{Z}}\in\mathbb{S}^{n}~|~{\mathbf{Z}}\succeq 0\}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then,
-$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{\textrm{rank}({\mathbf{X}})}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{\min}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$
-where $\lambda_{\min}({\mathbf{X}})>0$ denotes the smallest non-zero eigenvalue of ${\mathbf{X}}$.
+$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq 4\sqrt{\textrm{rank}({\mathbf{X}})}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
 Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\|{{\mathbf{X}}}\|_{1}\leq R,~\textrm{rank}({\mathbf{X}})\leq s\}$.
-In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=\left({2\sqrt{s}+\frac{2R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$ (note that for any feasible ${\mathbf{Y}}$ it holds that $\|{{\mathbf{Y}}}\|_{2}\leq R$).
-Here the hard-thresholding step in Algorithm 1 (computing the point $\widehat{{\mathbf{x}}}_{t}$) is done by computing the rank-$s$ approximation of the current iterate ${\mathbf{x}}_{t}$, which requires computing only the top $s$ components in the SVD of ${\mathbf{x}}_{t}$ (and discarding the rest).
+In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=16s$.
+[…]
 Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest signed eigenvalues) in the eigen-decomposition of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, ${\mathbf{u}}_{i}\in\mathbb{R}^{n},i=1,\dots,s$.
@@
 For the correctness of this procedure see Lemma 3 in the appendix.
-Overall, aside from the gradient computation, the runtime of each iteration of Algorithm 1 is dominated by computing two rank-$s$ eigen-decompositions of a $n\times n$ symmetric matrix, which is far more efficient than a full-rank eigen-decomposition once $s<<n$.
+[…]
 Corollary 2.
-When applied to Problem (9) with an optimal solution ${\mathbf{X}}^{*}$ which satisfies $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and $\textrm{Tr}({\mathbf{x}}^{*})=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\left({\sqrt{s}+\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}\frac{\beta_{1}}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of the gradient computations, the runtime of each iteration is dominated by two rank-$s$ eigen-decompositions of $n\times n$ symmetric matrices.
-Comparing the rate in Corollary 2 to the state-of-the-art non-Euclidean accelerated gradient method in [14], which has convergence rate $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$, we can see that our rate is not only unaccelerated (i.e., linear in the condition number $\frac{s\beta_{1}}{\alpha_{2}}$ as opposed to linear in its square-root), but we also have an additional term of $\left({\frac{R}{\lambda_{\min}({\mathbf{X}}^{*})}}\right)^{2}$, which may be significantly larger than 1.
-On the other-hand, the method in [14] requires a full eigen-decomposition of a $n\times n$ matrix each iteration, which is often prohibitive for large $n$.
+[…]
 Remark 2.
@@
 Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$.
-Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in $(0,1)$).
+Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in (0,1)(0,1)( 0 , 1 )).
 For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
@@
 5 Open Questions
-We conclude with several open questions which we believe are of interest.
+We conclude with two open questions which we believe are of interest.
 First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems.
-Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{s\beta_{1}}{\alpha_{2}}}$ while maintaining the sparse updates.
-Finally, when comparing the matrix Problem (9) with Problem (8), it is interesting if the additional term of $R^{2}/\lambda_{\min}({\mathbf{X}}^{*})^{2}$ in the condition number could be removed.
+Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
 Acknowledgments
@@
 Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$.
-Also, let ${\mathbf{Q}}\in\mathbb{S}^{n}$ denote the projection matrix onto subspace spanned by the top $k$ eigenvectors of ${\mathbf{Y}}$ (corresponding to the $k$ largest signed eigenvalues).
-Recall that $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm and Euclidean norm in $\mathbb{S}^{n}$, respectively.
-In the following we also let $\sigma_{i}({\mathbf{Z}})$ denote the $i$th largest singular value of some real matrix ${\mathbf{Z}}$.
-It holds that,
-$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp})({\mathbf{X}}-{\mathbf{Y}})}\|_{1}$
-$=\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})+{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
-$\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
-$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}}\|_{1}$
-$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{P}}{\mathbf{Y}}}\|_{1}$
-$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}-{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$=\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\underset{(a)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\underset{(b)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{X}}}\|_{1}-\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{1}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\leq\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\sqrt{k}\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}})}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\leq 2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$
-$\underset{(d)}{\leq}2\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}+\|{{\mathbf{Y}}}\|_{2}\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2},$
-where (a) follows from the assumption of the lemma, (b) follows since on one hand $\|{{\mathbf{Q}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, and on the other hand, using von Neumann’s inequality for the singular values (see for instance [19]), we have that $\|{{\mathbf{P}}{\mathbf{Y}}}\|_{1}=\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}}{\mathbf{Y}})\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{P}})\sigma_{i}({\mathbf{Y}})\leq\sum_{i=1}^{k}\sigma_{i}({\mathbf{Y}})$, where we have used the fact that $\textrm{rank}({\mathbf{P}})\leq k$ and $\|{{\mathbf{P}}}\|_{\infty}=1$.
-Inequality (c) follows since ${\mathbf{X}}={\mathbf{P}}{\mathbf{X}}$, and (d) follows again from von Neumann’s inequality and the Cauchy-Schwarz inequality:
-$\|{({\mathbf{Q}}-{\mathbf{P}}){\mathbf{Y}}}\|_{1}$ $\leq\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})\sigma_{i}({\mathbf{Y}})$
-$\leq\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Q}}-{\mathbf{P}})^{2}}\sqrt{\sum_{i=1}^{n}\sigma_{i}({\mathbf{Y}})^{2}}$
-$=\|{{\mathbf{Y}}}\|_{2}\cdot\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}.$
-Using the Davis-Kahan Theorem (see for instance Lemma A.1 in [11]) we have that,
-$\|{{\mathbf{Q}}-{\mathbf{P}}}\|_{2}\leq\frac{2\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}.$
-Thus, we conclude that
-$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq\left({2\sqrt{k}+\frac{2\|{{\mathbf{Y}}}\|_{2}}{\lambda_{k}({\mathbf{X}})}}\right)\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
+Throughout the proof we let $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm (sum of singular values) and the Euclidean (Frobenius) norm in some space of real (not necessarily symmetric) matrices, respectively.
+We begin by recalling two important facts.
+First, for any real (not necessarily symmetric) matrix ${\mathbf{A}}$ it holds that
+$\|{{\mathbf{A}}}\|_{1}\leq\sqrt{\textrm{rank}({\mathbf{A}})}\|{{\mathbf{A}}}\|_{2}.$ (9)
+Second, for any two real (not necessarily symmetric) matrices of appropriate dimension (such that the multiplication is defined) ${\mathbf{A}},{\mathbf{B}}$ it holds that
+$\|{{\mathbf{A}}{\mathbf{B}}}\|_{2}\leq\|{{\mathbf{A}}}\|_{\infty}\|{{\mathbf{B}}}\|_{2},$ (10)
+where we let $\|{{\mathbf{A}}}\|_{\infty}$ denote the largest singular value of ${\mathbf{A}}$.
+Let us denote ${\mathbf{D}}={\mathbf{X}}-{\mathbf{Y}}$. It holds that,
+$\|{{\mathbf{D}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp}){\mathbf{D}}({\mathbf{P}}+{\mathbf{P}}_{\perp})}\|_{1}$
+$\leq\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}+2\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}.$ (11)
+We continue by upper-bounding each of the three terms in the RHS separately.
+Note that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}})\leq k$. Thus, using Eq. (9) and then Eq. (10), and noting that $\|{{\mathbf{P}}}\|_{\infty}=1$, we have that
+$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (12)
+To bound the second term in the RHS of (B) we have,
+$\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}$ $=\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}}\|_{1}\underset{(a)}{=}\textrm{Tr}({\mathbf{P}}_{\perp}{\mathbf{Y}})=\textrm{Tr}(({\mathbf{I}}-{\mathbf{P}}){\mathbf{Y}})$
+$\underset{(b)}{\leq}\textrm{Tr}({\mathbf{X}})-\textrm{Tr}({\mathbf{P}}{\mathbf{Y}})=\textrm{Tr}({\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}})\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}}}\|_{1}$
+$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$ (13)
+where (a) holds since ${\mathbf{Y}}$ is positive semidefinite, and thus so is ${\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}$, (b) holds by the assumption that $\textrm{Tr}({\mathbf{Y}})=\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}=\textrm{Tr}({\mathbf{X}})$, and (c) is due to Eq.
+(12)
+Finally, for the last term in the RHS of (B), noting that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp})\leq k$, using Eq.
+(9) and then Eq.
+(10) (and noting again that $\|{{\mathbf{P}}}\|_{\infty}=\|{{\mathbf{P}}_{\perp}}\|_{\infty}=1$), we have that
+$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (14)
+The lemma follows from plugging-in (12), (B), and (14) into the RHS of (B). ∎
 Appendix C Sparse Projections onto the Bounded Positive Semidefinite Cone

### v3

@@
 Abstract
-It was recently established that for convex optimization problems with a sparse optimal solution (may it be entry-wise sparsity or matrix rank-wise sparsity) it is possible to have linear convergence rates which depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipchitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of the optimal solution.
-However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving also the runtime of each iteration, which may still be prohibitively high for high-dimensional problems.
+It was recently established that for convex optimization problems with sparse optimal solutions (be it entry-wise sparsity or matrix rank-wise sparsity) it is possible to design first-order methods with linear convergence rates that depend on an improved mixed-norm condition number of the form $\frac{\beta_{1}{}s}{\alpha_{2}}$, where $\beta_{1}$ is the $\ell_{1}$-Lipschitz continuity constant of the gradient, $\alpha_{2}$ is the $\ell_{2}$-quadratic growth constant, and $s$ is the sparsity of optimal solutions.
+However, beyond the improved convergence rate, these methods are unable to leverage the sparsity of optimal solutions towards improving the runtime of each iteration as well, which may still be prohibitively high for high-dimensional problems.
 In this work, we establish that linear convergence rates which depend on this improved condition number can be obtained using only sparse updates, which may result in overall significantly improved running times.
@@
 $\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (1)
-[…]
-Example 2: trace-constrained minimization for low-rank matrix recovery.
-$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}),$ (2)
-where $\mathbb{S}^{n}$ is the space of real symmetric $n\times{}n$ matrices and the relation ${\mathbf{X}}\succeq 0$ denotes ${\mathbf{X}}$ is positive semidefinite.
-Here we also assume $f$ is convex and continuously differentiable, and that Problem (2) indeed admits a unique optimal solution ${\mathbf{X}}^{*}$ such that $\textrm{rank}({\mathbf{X}}^{*})<<n$.
-[…]
-In recent works [16, 14], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that the optimal solution is indeed sparse/low-rank, it is possible to obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipchitz continuity constant of the gradient w.r.t.
+where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex and continuously differentiable, and we assume that all optimal solutions have sparsity (number of non-zero entries) at most $s<<n$.
+Example 2: Nuclear norm-constrained minimization for low-rank matrix recovery.
+$\min_{{\mathbf{X}}\in\mathbb{R}^{m\times n}:~\|{{\mathbf{X}}}\|_{1}\leq R}f({\mathbf{X}}),$ (2)
+where here we let $\|{\cdot}\|_{1}$ denote the matrix nuclear norm in $\mathbb{R}^{m\times n}$, i.e., sum of singular values.
+Here we also assume $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}$ is convex and continuously differentiable, and that all optimal solutions have rank at most $s<<\min\{m,n\}$.
+With regard to both problems, let us denote by $\beta$ the Lipschitz continuity constant of the gradient w.r.t.
+some norm $\|{\cdot}\|$, i.e., for all ${\mathbf{x}},{\mathbf{y}}$ we have that $\|{\nabla{}f({\mathbf{x}})-\nabla{}f({\mathbf{y}})}\|_{*}\leq\beta\|{{\mathbf{x}}-{\mathbf{y}}}\|$, where $\|{\cdot}\|_{*}$ denotes the norm dual to $\|{\cdot}\|$.
+Suppose also that $f$ satisfies a lower curvature condition such as strong convexity, or the weaker quadratic growth condition [22], with constant $\alpha>0$ w.r.t.
+the same norm, i.e., that for any feasible $\mathbf{x}$} it holds that $\min_{{\mathbf{x}}^{*}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{2}\leq\frac{2}{\alpha}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right)$, where we denote by $\mathcal{X}^{*}$ the set of all optimal solutions.
+In this setup, known gradient methods can produce an $\epsilon$-approximated solution (in function value) in worst-case number of iterations that scales like $\frac{\beta}{\alpha}\log{1/\epsilon}$, or even $\sqrt{\frac{\beta}{\alpha}}\log{1/\epsilon}$ when accelerated methods are used [24, 2, 22].
+The ratio $\beta/\alpha$ is the condition number of $f$ w.r.t.
+the norm $\|{\cdot}\|$ and is naturally dependent on the choice of norm.
+In recent works [17, 15], it was established that for convex relaxations of sparse optimization problems such as Problems (1) and (2), and under the quite plausible assumption that all optimal solutions are indeed sparse/low-rank, it is possible to design gradient methods that obtain linear convergence rates which depend on a mixed-norm condition number of the form $\frac{\beta_{1}s}{\alpha_{2}}$, where $\beta_{1}$ is the Lipschitz continuity constant of the gradient w.r.t.
 the $\ell_{1}$ norm, $\alpha_{2}$ is the quadratic growth constant w.r.t.
-the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of the optimal solution.
+the $\ell_{2}$ norm, and $s$ is an upper-bound on the sparsity of optimal solutions.
 This condition number could indeed be far better than the pure $\ell_{2}$ or pure $\ell_{1}$ condition numbers.
@@
 One can easily observe that the Euclidean condition number for this problem is given by $\frac{\beta_{2}}{\alpha_{2}}=\frac{\lambda_{\max}({\mathbf{A}})}{\lambda_{\min}({\mathbf{A}})}=\frac{n+\alpha}{\alpha}$.
-On the other-hand, a simple calculation shows that the $\ell_{1}$-Lipchitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$.
-To bound the $\ell_{1}$-quadratic growth constant we observe that due the the structure of the matrix ${\mathbf{A}}$,
+On the other hand, a simple calculation shows that the $\ell_{1}$-Lipschitz continuity constant of the gradient is given by $\beta_{1}=\max_{i,j}|{{\mathbf{A}}_{i,j}}|=1+\alpha$.
+To bound the $\ell_{1}$-quadratic growth constant we observe that due to the structure of the matrix $\mathbf{A}$},
 $\forall{\mathbf{x}}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\sum_{i=1}^{n}{\mathbf{x}}_{i}-\sum_{j=1}^{n}{\mathbf{x}}^{*}_{j}}\right)^{2}.$
-In particular, assuming ${\mathbf{x}}^{*}\geq 0$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., ${\mathbf{x}}^{*}$ lies on the boundary of the $\ell_{1}$ ball, as often expected) we have that
-$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~|~{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=R\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
-Thus, in worst-case, the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
-However, in the example above, if ${\mathbf{x}}^{*}$ satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
+In particular, assuming ${\mathbf{x}}^{*}\geq 0$ we have that
+$\forall{\mathbf{x}}\in\{{\mathbf{y}}\in\mathbb{R}^{n}~|~{\mathbf{y}}\geq 0,\|{{\mathbf{y}}}\|_{1}=\|{{\mathbf{x}}^{*}}\|_{1}\}:\quad f({\mathbf{x}})-f({\mathbf{x}}^{*})=\frac{\alpha}{2}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}.$
+This implies that the $\ell_{1}$-quadratic growth constant $\alpha_{1}$ is not better than $\frac{\alpha}{n}$, which implies that the $\ell_{1}$-condition number is $\frac{\beta_{1}}{\alpha_{1}}=\frac{n+n\alpha}{\alpha}$, which is no better than the Euclidean one.
+However, in the example above, if $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, the sparsity-based mixed-norm condition number satisfies $\frac{\beta_{1}s}{\alpha_{2}}=\frac{s(1+\alpha)}{\alpha}$, which is in particular independent of the ambient dimension, and thus may lead to far-better (worst-case) convergence rates for suitable first-order methods.
 The example in Eq.
-(3) could be directly extended also to Problem (8) by taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\textrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$, where $\|{\cdot}\|_{2}$ here is the Euclidean norm in $\mathbb{S}^{n}$, and instead of the entry-wise $\ell_{1}$ norm we consider the Schatten 1-norm (i.e., the $\ell_{1}$ norm w.r.t.
-the eigenvalues), and $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$.
-The recent works [16, 14]111while [16, 14] consider a stochastic convex optimization setting in which the objective function $f$ is given as an expectation w.r.t.
-unknown distribution from which we can sample, here our focus is on the standard deterministic first-order convex optimization model indeed lead to convergence rates of the form $\tilde{O}\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ [16] and even an accelerated rate of the form $\tilde{O}\left({\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}\log\frac{1}{\epsilon}}\right)$ [14] (the $\tilde{O}(\cdot)$ notation hides certain poly-logarithmic factors), by applying a restarted non-Euclidean gradient method222these methods rely on Bregman proximal gradient computations with the Bregman distance induced by a squared $\ell_{p}$ norm for an appropriate choice of $p$ that best leverages the strong convexity w.r.t.
-the $\ell_{1}$ norm or its accelerated variant, together with hard-thresholding steps to sparsify intermediate iterates, see for instance Corollary 5 in [14].
-While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is large.
-For instance, in the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next.
-In the case of Problem (2) the time required by these methods to project onto the positive semidefinite cone ($O(n^{3})$ in standard implementations) may be prohibitive.
-In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the improved condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates which depend only on its square root).
-The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solution), which often allows for significant improvements in the runtime of each iteration.
+(3) could be directly extended also to Problem (2) by considering the space $\mathbb{R}^{n\times n}$ and taking $f({\mathbf{X}})=\frac{\alpha}{2}\|{{\mathbf{X}}-{\mathbf{X}}^{*}}\|_{2}^{2}+\frac{1}{2}\left({\mathrm{Tr}({\mathbf{X}}-{\mathbf{X}}^{*})}\right)^{2}$ with $\mathbf{X}$$^{*}$ satisfying $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<n$, and letting $\|{\cdot}\|_{2}$ here denote the Euclidean (Frobenius) norm, and instead of the entry-wise $\ell_{1}$ norm, we consider the matrix nuclear norm ($\ell_{1}$ norm w.r.t.
+the vector of singular values).
+[…]
+While these non-Euclidean gradient methods lead to superior convergence rates, they do not leverage the sparsity of optimal solutions towards additional computational gains, and in particular may still require prohibitive running times per iteration when the dimension is very high.
+In the case of Problem (1), these may suffer from prohibitive running time for updating the gradient direction from one iteration to the next.
+In the case of Problem (2), the time required by these methods to project onto the nuclear norm ball ($O(\min\{m,n\}^{2}\max\{m,n\})$ time, in standard implementations, to compute a full compact-form SVD) may be prohibitive.
+In this work we present a new and simple approach towards obtaining linear convergence rates that depend on the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ (though our method only obtains rates that are linear in it and not accelerated rates that scale with its square root).
+The main novel aspect is that our algorithm only applies sparse updates (with the same level of sparsity assumed for the optimal solutions), which often allows for significant improvements in the runtime of each iteration.
 In particular, our algorithm updates its iterates similarly to the classical Frank-Wolfe algorithm [16]:
-${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}\quad\textrm{for some}\quad\eta\in[0,1],$ (4)
-where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for ${\mathbf{x}}^{*}$.
+${\mathbf{x}}_{t+1}\leftarrow(1-\gamma){\mathbf{x}}_{t}+\gamma{\mathbf{v}}_{t}\quad\textrm{for some}\quad\gamma\in[0,1],$ (4)
+where ${\mathbf{v}}_{t}$ is a feasible point with the same level of sparsity assumed for the optimal solutions.
 For instance, in case of Problem (1) with a quadratic objective function $f({\mathbf{x}})={\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the fact that ${\mathbf{v}}_{t}$ is $s$-sparse, reduces the time to update the gradient direction from $O(n^{2})$ to $O(sn)$.
-In the case of Problem (2), since $\textrm{rank}({\mathbf{v}}_{t})\leq s$, computing it won’t require a full eigen-decomposition ($O(n^{3})$ time in standard implementations), but only computing the $s$ leading components in the eigen-decomposition, which is far more efficient.
+In the case of Problem (2), the fact that $\mathrm{rank}({\mathbf{v}}_{t})\leq s$ implies that computing it will not require a full compact-form SVD ($O(\min\{m,n\}^{2}\max\{m,n\})$ time in standard implementations), but only computing the $s$ leading components in the SVD, which is far more efficient.
 Two additional advantages which make our approach considerably simpler (compared to [17, 15]) are: I.
@@
 the updates in our algorithm (the way ${\mathbf{v}}_{t}$ in Eq.
-(4) is computed) are Euclidean, which often leads to significantly simpler computations.
-Compared to [16, 14], we do have an additional mild assumption, which in the case of Problems (1) and (2) above, requires that the optimal solution is on the boundary in the sense that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ or $\textrm{Tr}({\mathbf{X}}^{*})=R$, respectively.
-We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm regularization is redundant).
+(4) is computed) are Euclidean, which often leads to significantly simpler implementations.
+Compared to [17, 15], we do have an additional mild assumption, which in the case of Problems (1) and (2) requires that all optimal solutions are on the boundary, i.e., that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ for any optimal solution $\mathbf{x}$$^{*}$.
+We consider this to be quite mild since in the context of sparse recovery it is not likely that the optimal solution lies in the interior (in particular it implies that the norm constraint is redundant).
+Aside from our original method we also make the observation that for Problem (1), another method that is able to attain linear convergence with the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ that uses only sparse updates is the well-known away-steps Frank-Wolfe (AFW) method [14, 20].
+In Section A we show that, assuming all optimal solutions to Problem (1) are $s$-sparse and lie on the boundary (same assumption made for our proposed method), AFW has linear convergence of the form $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$.
+Note the additional dependence on the squared radius $R^{2}$.
+In particular, under the plausible assumption that $R=\Omega(s)$ (e.g., entries in optimal solutions are $\Theta(1)$), this rate becomes $\Omega\left({\frac{\beta_{1}s^{3}}{\alpha_{2}}\log{1/\epsilon}}\right)$.
+Note also that AFW does not extend to handle Problem (2).
+Algorithm required parameters #iterations to $\epsilon$ error iteration time
+minimization over $\ell_{1}$-ball with $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, ${\mathbf{A}}\succeq 0$, $\mathrm{nnz}({\mathbf{x}}^{*})\leq s~\forall{\mathbf{x}}^{*}\in\mathcal{X}^{*}$
+Away-Step Frank-Wolfe with line-search [14, 20] (see Section A) - $\frac{\beta_{1}sR^{2}}{\alpha_{2}}$ $O(n)$
+Restarted Acc. Grad. with hard-thresholding [15] (see Section B) $s,\beta_{1},\alpha_{2}$ $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ $O(n^{2})$
+Our Algorithm 1 (see Corollary 1) $s,\alpha_{2}$ $\frac{\beta_{1}s}{\alpha_{2}}$ $O(sn)$
+minimization over matrix nuclear norm ball, $\mathrm{rank}({\mathbf{X}}^{*})\leq s~\forall{\mathbf{X}}^{*}\in\mathcal{X}^{*}$
+Restarted Acc. Grad. with hard-thresholding [15] (see Section B) $s,\beta_{1},\alpha_{2}$ $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ full SVD of $m\times n$ matrix
+Our Algorithm 1 (see Corollary 2) $s,\alpha_{2}$ $\frac{\beta_{1}s}{\alpha_{2}}$ rank-$s$ SVD of $m\times n$ matrix
+Table 1: Comparison of first-order methods with linear convergence rates that scale with the mixed-norm condition number $\frac{\beta_{1}s}{\alpha_{2}}$ for Problems (1) and (2), omitting logarithmic factors and universal constants.
+From Table 1 we can see for instance that in case of Problem (1) with a quadratic objective, whenever $s<<\min\{R^{2},n^{2/3}\left({\alpha_{2}/\beta_{1}}\right)^{1/3}\}$, the overall runtime of our method is the state-of-the-art.
 The rest of the paper is organized as follows.
-In Section 2 we present a unified setup and optimization problem which captures the spare convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis.
+In Section 2 we present a unified setup and optimization problem which captures the sparse convex optimization problems we are interested in, a generic algorithm for solving this problem, and its convergence analysis.
 In Section 3 we instantiate our generic algorithm in the special cases of Problems (1) and (2).
@@
 $\min_{{\mathbf{x}}\in\mathcal{K}}f({\mathbf{x}}),$ (5)
-[…]
+where $\mathcal{K}$ is a convex and compact subset of a Euclidean vector space $\mathbb{E}$ equipped with an inner-product $\langle{\cdot,\cdot}\rangle$, $f$ is convex over $\mathcal{K}$ with Lipschitz continuous gradient with parameter $\beta$ w.r.t.
+some norm $\|{\cdot}\|$, i.e.,
 $\forall{\mathbf{x}},{\mathbf{y}}\in\mathcal{K}:\quad\|{\nabla{}f({\mathbf{x}})-\nabla{}f({\mathbf{y}})}\|_{*}\leq\beta\|{{\mathbf{x}}-{\mathbf{y}}}\|,$
 where $\|{\cdot}\|_{*}$ denotes the dual norm.
+We let $f^{*}$ denote the optimal value of (5) and we let $\mathcal{X}^{*}\subseteq\mathcal{K}$ denote the set of minimizers.
 We let $\|{{\mathbf{x}}}\|_{2}=\sqrt{\langle{{\mathbf{x}},{\mathbf{x}}}\rangle}$ denote the Euclidean norm.
-We shall assume Problem (5) admits a unique optimal solution ${\mathbf{x}}^{*}$, and satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
-$\forall{\mathbf{x}}\in\mathcal{K}:\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (6)
+We shall assume Problem (5) satisfies a quadratic growth condition with parameter $\alpha_{2}>0$ w.r.t. the Euclidean norm, i.e.,
+$\forall{\mathbf{x}}\in\mathcal{K}:\quad\min_{{\mathbf{x}}^{*}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}^{2}\leq\frac{2}{\alpha_{2}}\left({f({\mathbf{x}})-f({\mathbf{x}}^{*})}\right).$ (6)
 We now make an additional assumption that further specializes Problem (5).
-In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for the optimal solution ${\mathbf{x}}^{*}$.
+In Section 3 we show that in the case of Problems (1) and (2) (and certain generalizations), Assumption 1 is satisfied once we make the plausible assumption that the norm constraint is tight for all optimal solutions.
 Assumption 1.
-There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such ${\mathbf{x}}^{*}\in\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
+There exists a closed subset $\mathbb{E}_{s}\subset\mathbb{E}$ such that $\mathcal{X}^{*}\subseteq\mathcal{K}_{s}:=\mathcal{K}\cap\mathbb{E}_{s}$, and there exist scalars $s,s_{\mathcal{K}}^{*}\geq 1$ such that,
   1. 1.
-    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
+    for all ${\mathbf{x}},{\mathbf{y}}\in\mathbb{E}_{s}$: $\|{{\mathbf{x}}-{\mathbf{y}}}\|\leq\sqrt{2s}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}$,
   2. 2.
-    for all ${\mathbf{x}}\in\mathcal{K}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
+    for all ${\mathbf{x}}\in\mathcal{K}$ and ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$: $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|\leq\sqrt{s^{*}_{\mathcal{K}}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|_{2}$,
   3. 3.
-    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathcal{K}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
-In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors which are $s$-sparse (in some sense), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to $s$-sparse vectors.
+    for all ${\mathbf{x}}\in\mathcal{K}$: $\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|=\arg\min\nolimits_{{\mathbf{y}}\in\mathbb{E}_{s}}\|{{\mathbf{y}}-{\mathbf{x}}}\|_{2}$ (note the argmin need not be a singleton),
+In the context of Problems (1) and (2) we should think of $\mathbb{E}_{s}$ as the restriction of the space $\mathbb{E}$ to vectors with a bounded level of sparsity (either in the sense of number of non-zero entries or matrix rank), and accordingly $\mathcal{K}_{s}$ is the restriction of $\mathcal{K}$ to vectors in $\mathbb{E}_{s}$.
 Additionally, while not part of Assumption 1, we shall implicitly assume that it is computationally efficient to compute a Euclidean projection onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$ (in case the projection is not a singleton, compute one element in the set of possible projections).
-Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively.
+Accordingly, we denote by $\Pi_{\mathbb{E}_{s}}[\cdot]$ and $\Pi_{\mathcal{K}_{s}}[\cdot]$ the Euclidean projection operation onto $\mathbb{E}_{s}$ and $\mathcal{K}_{s}$, respectively.
+Note that due to Item 3 in Assumption 1, $\Pi_{\mathbb{E}_{s}}[\cdot]$ is also the projection w.r.t.
+the norm $\|{\cdot}\|$.
+In the context of both Problems (1) and (2), we shall demonstrate in the sequel that computing these projections is much more efficient than computing the Euclidean projection over the entire feasible set $\mathcal{K}$ when the optimal solutions are indeed sparse.
 We can now present our algorithm for solving the generic Problem (5) (under the assumptions listed above).
 Our algorithm is composed of three simple ingredients.
-First, we use a Frank-Wolfe like scheme [15] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq.
+First, we use a Frank-Wolfe-like scheme [16] in which the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ are updated by taking a convex combination between the current feasible iterate and a new feasible point, as already shown in Eq.
 (4).
 Second, and differently from the classical Frank-Wolfe method in which the new point ${\mathbf{v}}_{t}$ is the output of a linear optimization oracle, here we take it to be a Euclidean proximal gradient step w.r.t.
-to restricted set $\mathcal{K}_{s}$ (instead of w.r.t.
+restricted set $\mathcal{K}_{s}$ (instead of w.r.t.
 the original set $\mathcal{K}$).
-This is similar to approach taken in previous works [1, 10, 12, 13] (also referred to as a weak proximal oracle [10, 13]).
+This is similar to the approach taken in previous works [1, 11, 12, 13] (also referred to as a weak proximal oracle [11, 13]).
 Third, the $\ell_{2}$-term in this proximal gradient step is not w.r.t.
-to the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
+the current iterate ${\mathbf{x}}_{t}$, but with respect to its projection onto $\mathbb{E}_{s}$, which should be understood as a hard-thresholding (or sparsification) of ${\mathbf{x}}_{t}$.
 Algorithm 1 Frank-Wolfe with sparse proximal oracle and hard-thresholding
@@
   $\widehat{{\mathbf{x}}}_{t}\leftarrow\Pi_{\mathbb{E}_{s}}[{\mathbf{x}}_{t}]$
-  ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{2s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
-  ${\mathbf{x}}_{t+1}\leftarrow(1-\eta){\mathbf{x}}_{t}+\eta{\mathbf{v}}_{t}$
+  ${\mathbf{v}}_{t}\leftarrow\arg\min\nolimits_{{\mathbf{v}}\in\mathcal{K}_{s}}\langle{{\mathbf{v}}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2s\beta\eta\|{{\mathbf{v}}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}$ {equivalently, ${\mathbf{v}}_{t}\leftarrow\Pi_{\mathcal{K}_{s}}[\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})]$}
+  set $\gamma_{t}\in[0,1]$ either by line-search: $\gamma_{t}\leftarrow\arg\min_{\gamma\in[0,1]}f((1-\gamma){\mathbf{x}}_{t}+\gamma{\mathbf{v}}_{t})$ OR use a pre-defined $\gamma_{t}\in[0,1]$
+  ${\mathbf{x}}_{t+1}\leftarrow(1-\gamma_{t}){\mathbf{x}}_{t}+\gamma_{t}{\mathbf{v}}_{t}$
 end for
 Theorem 1.
-Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with step-size $\eta=\frac{\alpha_{2}}{4\beta(4s+s^{*}_{\mathcal{K}})}$. Then,
-$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(4s+s^{*}_{\mathcal{K}})}}\right).$
-Proof.
-Fix some iteration $t$ of the algorithm. The Lipchitz continuity of the gradient implies the following well known inequality:
+Consider the iterates $({\mathbf{x}}_{t})_{t\geq 1}$ of Algorithm 1 with a fixed step-size $\eta=\frac{\alpha_{2}}{4\beta(8s+s^{*}_{\mathcal{K}})}\in[0,1]$ and when $\gamma_{t}$ is set either via line-search or to the pre-defined value $\gamma_{t}=\eta$ for all $t$.
+Then,
+$\forall t\geq 1:\quad f({\mathbf{x}}_{t+1})-f^{*}\leq\left({f({\mathbf{x}}_{t})-f^{*}}\right)\left({1-\frac{\alpha_{2}}{8\beta(8s+s^{*}_{\mathcal{K}})}}\right).$
+Before we prove the theorem let us make an important remark.
+Remark 1.
+Note that with the value of $\eta$ listed in the theorem, the update for ${\mathbf{v}}_{t}$ depends on the coefficient $s\beta\eta=\frac{s\alpha_{2}}{4(8s+s_{\mathcal{K}}^{*})}$.
+First, this is independent of $\beta$.
+Second, as we shall detail in Section 3, in both the context of Problem (1) and Problem (2), we shall have $s_{\mathcal{K}}^{*}=\Theta(s)$.
+Thus, aside from the sparsity level $s$, this coefficient depends on a single parameter which is the quadratic growth constant $\alpha_{2}$.
+While setting this constant can be difficult in general, since Algorithm 1 is a descent method (as the proof will establish) there is a simple practical approach to circumvent this difficulty: on each iteration, try a few possible values, compute ${\mathbf{v}}_{t}$ for each possible value and take the one which reduces the function value the most (in particular such computations can run in parallel).
+Proof of Theorem 1.
+Fix some iteration $t$ of the algorithm.
+Whether $\gamma_{t}$ is chosen via line-search or using the pre-defined value $\eta$ stated in the theorem, the Lipschitz continuity of the gradient implies the following well known inequality:
 $f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\frac{\beta\eta^{2}}{2}\|{{\mathbf{v}}_{t}-{\mathbf{x}}_{t}}\|^{2}$
-$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
+$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2},$
+where the second inequality is due to the triangle inequality and the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$.
 Since $\widehat{{\mathbf{x}}}_{t}$ and ${\mathbf{v}}_{t}$ are in $\mathbb{E}_{s}$, using Item 1 in Assumption 1 we have that,
-$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
-Using the definition of ${\mathbf{v}}_{t}$ in the algorithm and the assumption that ${\mathbf{x}}^{*}\in\mathcal{K}_{s}$ we have that,
-$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
-Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm, Item 3 and then Item 2 in Assumption 1 we have that,
-$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
+$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{v}}_{t}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{v}}_{t}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2},$
+In the following let ${\mathbf{x}}_{t}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}^{*}}\|{{\mathbf{x}}-{\mathbf{x}}_{t}}\|_{2}$. Using the definition of ${\mathbf{v}}_{t}$ in the algorithm we have that
+$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{x}}_{t}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}.$
+Using the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm with the assumption that ${\mathbf{x}}_{t}^{*}\in\mathcal{K}_{s}\subseteq\mathbb{E}_{s}$, and then Item 2 in Assumption 1 we have that,
+$\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|^{2}\leq\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|^{2}\leq s_{\mathcal{K}}^{*}\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}.$
 Combining the last two inequalities we have that,
-$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
-$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(2s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+2\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
-$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
-where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$ in the algorithm.
+$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+2\beta\eta^{2}s\|{{\mathbf{x}}_{t}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{2}^{2}+\beta\eta^{2}s^{*}_{\mathcal{K}}\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}$
+$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2}+4\beta\eta^{2}s\|{\widehat{{\mathbf{x}}}_{t}-{\mathbf{x}}_{t}}\|_{2}^{2}$
+$\leq f({\mathbf{x}}_{t})+\eta\langle{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t},\nabla{}f({\mathbf{x}}_{t})}\rangle+\beta\eta^{2}(8s+s^{*}_{\mathcal{K}})\|{{\mathbf{x}}_{t}^{*}-{\mathbf{x}}_{t}}\|_{2}^{2},$
+where the last inequality follows from the definition of $\widehat{{\mathbf{x}}}_{t}$.
 Using both the convexity and quadratic growth of $f$ (Eq. (6)), we have that
-$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right)+\frac{2\beta\eta^{2}(4s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f({\mathbf{x}}^{*})}\right).$
-Subtracting $f({\mathbf{x}}^{*})$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
+$f({\mathbf{x}}_{t+1})$ $\leq f({\mathbf{x}}_{t})-\eta\left({f({\mathbf{x}}_{t})-f^{*}}\right)+\frac{2\beta\eta^{2}(8s+s^{*}_{\mathcal{K}})}{\alpha_{2}}\left({f({\mathbf{x}}_{t})-f^{*}}\right).$
+Subtracting $f^{*}$ from both sides and plugging-in our choice of step-size $\eta$, completes the proof.
 3 Applications
@@
 $\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\|{{\mathbf{x}}}\|_{1}\leq R}f({\mathbf{x}}),$ (7)
-[…]
-We shall require the following simple lemma (a proof is brought in the appendix for completeness).
+Here we instantiate our Algorithm 1 by setting $\|{\cdot}\|$ to be the $\ell_{1}$ norm $\|{\cdot}\|_{1}$ in $\mathbb{R}^{n}$ (and accordingly $\|{\cdot}\|_{*}=\|{\cdot}\|_{\infty}$).
+As already discussed in the Introduction, we assume there is a positive integer $s<<n$ such that every optimal solution $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s$ and also $\|{{\mathbf{x}}^{*}}\|_{1}=R$ (i.e., all optimal solutions lie on a low-dimensional face of the $\ell_{1}$ ball).
+We shall require the following simple lemma (a proof is given in the appendix for completeness).
 Lemma 1.
 Let ${\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{n}$ such that $\|{{\mathbf{y}}}\|_{1}\leq\|{{\mathbf{x}}}\|_{1}$. Then,
-$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\textrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
-Let $s$ be such that $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\textrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\|{{\mathbf{x}}}\|_{1}\leq R,~\textrm{nnz}({\mathbf{x}})\leq s\}$.
-In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=2s$.
+$\|{{\mathbf{x}}-{\mathbf{y}}}\|_{1}\leq 2\sqrt{\mathrm{nnz}({\mathbf{x}})}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
+Let $s$ be such that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\mathrm{nnz}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{x}}\in\mathbb{R}^{n}~|~\|{{\mathbf{x}}}\|_{1}\leq R,~\mathrm{nnz}({\mathbf{x}})\leq s\}$, i.e., $\mathcal{K}_{s}$ is the restriction of the $\ell_{1}$ ball to $s$-sparse vectors.
+In light of Lemma 1 and the assumption that $\|{{\mathbf{x}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=4s$.
 Note that computing the hard-thresholding step in Algorithm 1 (the point $\widehat{{\mathbf{x}}}_{t}$) is straightforward by zeroing-out all but the largest (in absolute value) entries in ${\mathbf{x}}_{t}$.
-Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$ onto the $s$-sparsified $\ell_{1}$-ball $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the $n-s$ lower entries in ${\mathbf{z}}_{t}$) to obtain again a point the $n$-dimensional $\ell_{1}$-ball of radius $R$.
-See [3] for proof of the correctness of this procedure.
+Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, by computing the Euclidean projection of the point ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})$ onto $\mathcal{K}_{s}$, is done by computing the Euclidean projection of the vector in $\mathbb{R}^{s}$ which corresponds to the $s$-largest (in absolute value) entries in ${\mathbf{z}}_{t}$ onto the $s$-dimensional $\ell_{1}$-ball of radius $R$, and then appending zeros (in the entries which correspond to the nsn-s lower entries in ${\mathbf{z}}_{t}$) to obtain again a point in the $n$-dimensional $\ell_{1}$-ball of radius $R$.
+See [3] for a proof of the correctness of this procedure.
 Overall, aside from the gradient computation, each iteration of Algorithm 1 requires $O(n\log{}s)$ time (which is the time to find the $s$ largest entries in an array of length $n$ — the most time consuming operation).
 Corollary 1.
-When applied to Problem (7) with an optimal solution ${\mathbf{x}}^{*}$ which satisfies $\textrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside of gradient computations, the runtime of each iteration is $O(n\log{s})$.
-In particular, if $f$ is (convex) quadratic, the overall runtime of each iteration is $O(sn)$.
+When applied to Problem (7), and assuming any optimal solution $\mathbf{x}$$^{*}$ satisfies $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside from gradient computations, the runtime of each iteration is $O(n\log{s})$.
+In particular, if $f$ is (convex) quadratic of the form $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$, the overall runtime of each iteration is $O(sn)$.
 Remark 2.
-It is straightforward to extend the result above to the case in which ${\mathbf{x}}$ is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~|~\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as ${\mathbf{x}}^{*}$ satisfies $\|{{\mathbf{x}}}\|_{1}=R$.
+It is straightforward to extend the result above to the case in which $\mathbf{x}$} is constrained to be in $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{x}}~|~\|{{\mathbf{x}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{n}$ is closed and convex, as long as every optimal solution $\mathbf{x}$$^{*}$ satisfies $\|{{\mathbf{x}}^{*}}\|_{1}=R$.
 For instance when $\mathcal{C}=\{{\mathbf{x}}~|~{\mathbf{x}}\geq 0,\sum_{i=1}^{n}{\mathbf{x}}_{i}=R\}$, we have that $\mathcal{K}$ is simply a scaled simplex for which the assumption $\|{{\mathbf{x}}^{*}}\|_{1}=R$ always holds.
 Moreover, when $\mathcal{C}$ is closed under permutation of coordinates (e.g., an $\ell_{p}$ ball or its non-negative restriction), the sparse proximal step (computation of ${\mathbf{v}}_{t}$ in Algorithm 1) admits efficient implementation as detailed in [3].
-3.2 Bounded positive semidefinite cone setup
+3.2 Nuclear norm ball setup
 We recall the second problem in our Introduction:
-$\min_{{\mathbf{X}}\in\mathbb{S}^{n}:~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R}f({\mathbf{X}}).$ (8)
-[…]
-The following lemma is analogues to Lemma 1. A proof is given in the appendix for completeness.
+$\min_{{\mathbf{X}}\in\mathbb{R}^{m\times n}:~\|{{\mathbf{X}}}\|_{1}\leq R}f({\mathbf{X}}).$ (8)
+Here we instantiate our Algorithm 1 by setting $\|{\cdot}\|$ to be the matrix nuclear norm $\|{{\mathbf{X}}}\|_{1}=\sum_{i=1}^{\min\{m,n\}}|{\sigma_{i}({\mathbf{X}})}|$, and accordingly $\|{{\mathbf{X}}}\|_{*}=\|{{\mathbf{X}}}\|_{\mathrm{op}}=\max_{i\in[\min\{m,n\}]}\sigma_{i}({\mathbf{X}})$ which is the operator norm (aka the spectral norm).
+Here also we let $\|{\cdot}\|_{2}$ denote the Euclidean (Frobenius) matrix norm.
+As already discussed in the Introduction, we assume there is a positive integer $s<<\min\{m,n\}$ such that every optimal solution $\mathbf{X}$$^{*}$ satisfies $\mathrm{rank}({\mathbf{X}}^{*})\leq s$ and also $\|{{\mathbf{X}}^{*}}\|_{1}=R$ (i.e., all optimal solutions lie on a low-dimensional face of the nuclear norm ball).
+The following lemma is analogous to Lemma 1.
+A proof is given in the appendix for completeness.
 Lemma 2.
-Let ${\mathbf{X}},{\mathbf{Y}}\in\{{\mathbf{Z}}\in\mathbb{S}^{n}~|~{\mathbf{Z}}\succeq 0\}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then,
-$\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{1}\leq 4\sqrt{\textrm{rank}({\mathbf{X}})}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2}.$
-Let $s$ be such that $\textrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~\textrm{rank}({\mathbf{x}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\|{{\mathbf{X}}}\|_{1}\leq R,~\textrm{rank}({\mathbf{X}})\leq s\}$.
-In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=16s$.
-[…]
-Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest signed eigenvalues) in the eigen-decomposition of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta^{2}}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, ${\mathbf{u}}_{i}\in\mathbb{R}^{n},i=1,\dots,s$.
-Then, project the vector of eigenvalues $(\lambda_{1},\dots,\lambda_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~|~{\mathbf{y}}\geq 0,~\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of eigenvalues $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$.
-Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ (i.e., the eigenvectors remain unchanged).
+Let ${\mathbf{X}},{\mathbf{Y}}\in\mathbb{R}^{m\times n}$ such that ${\mathbf{X}}\neq\mathbf{0}$, $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$. Then
+$\|{\mathbf{X}}-{\mathbf{Y}}\|_{1}\leq 2\sqrt{2\mathrm{rank}({\mathbf{X}})}\,\|{\mathbf{X}}-{\mathbf{Y}}\|_{2}.$
+Let $s$ be such that $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<n$ and let us set $\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\mathrm{rank}({\mathbf{X}})\leq s\}$ which implies that $\mathcal{K}_{s}=\mathcal{K}\cap\mathbb{E}_{s}=\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,~\mathrm{rank}({\mathbf{X}})\leq s\}$, i.e., $\mathcal{K}_{s}$ is the restriction of nuclear norm ball to matrices with rank at most $s$.
+In light of Lemma 2 and the assumption that $\|{{\mathbf{X}}^{*}}\|_{1}=R$ it is straightforward to verify that the conditions in Assumption 1 indeed hold with constants $s$ and $s^{*}_{\mathcal{K}}=8s$.
+Here the hard-thresholding step in Algorithm 1 (computing the point $\widehat{{\mathbf{x}}}_{t}$) is done by computing the rank-$s$ approximation of the current iterate ${\mathbf{x}}_{t}$, which requires computing only the top $s$ components in the SVD of ${\mathbf{x}}_{t}$ (and discarding the rest).
+Computing the sparse proximal step, the point ${\mathbf{v}}_{t}$, is done as follows: compute the $s$ top components (corresponding to the largest singular values) in the SVD of the matrix ${\mathbf{z}}_{t}=\widehat{{\mathbf{x}}}_{t}-\frac{1}{4s\beta\eta}\nabla{}f({\mathbf{x}}_{t})$, and denote it as $\sum_{i=1}^{s}\sigma_{i}{\mathbf{u}}^{(i)}{\mathbf{v}}^{(i)\top}$, ${\mathbf{u}}^{(i)}\in\mathbb{R}^{m},{\mathbf{v}}^{(i)}\in\mathbb{R}^{n},i=1,\dots,s$.
+Then, project the vector of singular values $(\sigma_{1},\dots,\sigma_{s})$ onto the $s$-dimensional downward closed simplex $\mathcal{S}_{R}:=\{{\mathbf{y}}\in\mathbb{R}^{s}~|~{\mathbf{y}}\geq 0,~\sum_{i=1}^{s}{\mathbf{y}}_{i}\leq R\}$ to obtain the new vector of singular values $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$.
+Then, we set ${\mathbf{v}}_{t}=\sum_{i=1}^{s}\sigma_{i}^{\prime}{\mathbf{u}}^{(i)}{\mathbf{v}}^{(i)\top}$ (i.e., the singular vectors remain unchanged).
 For the correctness of this procedure see Lemma 4 in the appendix.
-[…]
+Overall, aside from the gradient computation, the runtime of each iteration of Algorithm 1 is dominated by computing two rank-$s$ SVDs of a $m\times n$ matrix, which is far more efficient than a full-rank SVD when $s<<\min\{m,n\}$.
 Corollary 2.
-[…]
+When applied to Problem (8), and assuming any optimal solution $\mathbf{X}$$^{*}$ satisfies $\mathrm{rank}({\mathbf{X}}^{*})\leq s<<\min\{m,n\}$ and $\|{{\mathbf{X}}^{*}}\|_{1}=R$, Algorithm 1 admits an implementation that finds an $\epsilon$-approximated solution (in function value) in $O\left({\frac{\beta_{1}s}{\alpha_{2}}\log\frac{1}{\epsilon}}\right)$ iterations, and aside from the gradient computations, the runtime of each iteration is dominated by two rank-$s$ SVDs of $m\times n$ matrices.
 Remark 3.
-As in Remark 1, here also we can extend Problem (8) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R\}$, where $\mathcal{C}\subset\mathbb{S}^{n}$ is closed and convex, as long as $\textrm{Tr}({\mathbf{X}}^{*})=R$.
-In particular, when $\mathcal{C}$ is a ball w.r.t.
-a Schatten $p$-norm ($\ell_{p}$ norm w.r.t.
-the eigenvalues), one can extend Lemma 3 accordingly.
-4 Some Numerical Evidence
-We present simple numerical simulations to demonstrate the benefit of the improved convergence rate of our method, which depends on a mixed-norm condition number, in the context of the $\ell_{1}$-constrained Problem (1).
+As in Remark 2, here also we can extend Problem (8) to include slightly more involved constraints of the form $\mathcal{K}=\mathcal{C}\cap\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R\}$, where $\mathcal{C}\subset\mathbb{R}^{m\times n}$ is closed and convex, as long as $\|{{\mathbf{X}}^{*}}\|_{1}=R$ for any optimal solution $\mathbf{X}$$^{*}$.
+For instance, setting $m=n$ and $\mathcal{C}=\mathbb{S}^{n}_{+}$ — the positive semidefinite (PSD) cone, we can obtain the analogue of Problem (8) for optimization over the trace-bounded PSD cone.
+In particular, one can extend Lemma 4 (efficient sparse prox computations) accordingly.
+4 Numerical Evidence
+We present simple numerical simulations on the $\ell_{1}$-constrained Problem (1).
 Similarly to the example in Eq.
 (3), we set $f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}({\mathbf{I}}+3\cdot\mathbf{1}\mathbf{1}^{\top})({\mathbf{x}}-{\mathbf{x}}^{*})$, where $\mathbf{1}$ denotes the all-ones vector.
-For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipchitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively.
-We compare the convergence rate of our Algorithm 1 to a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]).
-We set the dimension $n=3000$, the radius of the $\ell_{1}$ ball to $R=10$.
-The optimal solution ${\mathbf{x}}^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to ±$\pm\frac{R}{\textrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\textrm{nnz}({\mathbf{x}}^{*})$ is set in advanced, and for each level of sparsity we run 10 i.i.d.
-experiments (where in each one we resample ${\mathbf{x}}^{*}$).
-Accordingly, the figure below presents the average approximation error over 10 i.i.d.
-experiments.
-When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$.
-Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$.
-Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in (0,1)(0,1)( 0 , 1 )).
+For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipschitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively.
+In all experiments we set the dimension to $n=1000$, the radius of the $\ell_{1}$ ball to $R=10$.
+The optimal solution $\mathbf{x}$$^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to $\pm\frac{R}{\mathrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\mathrm{nnz}({\mathbf{x}}^{*})$ is set in advance, and for each level of sparsity we run 10 i.i.d.
+experiments (where in each one we resample $\mathbf{x}$$^{*}$) and display the average approximation error vs.
+number of iterations in Figure 1.
+Aside from our Algorithm 1, we implemented three benchmarks: I.
+the away-steps Frank-Wolfe algorithm with line-search that for this setting has worst-case convergence rate $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$ (see Section A), II.
+a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]), and III.
+a restarted non-Euclidean accelerated gradient method with hard-thresholding steps as described in Section B which has convergence rate $\tilde{O}\left({\sqrt{\frac{\beta_{1}{}s\ln(n)}{\alpha_{2}}}\log{1/\epsilon}}\right)$ (per Remark 4 we implemented Nesterov’s Method 5.6 [23] as our non-Euclidean accelerated gradient method).
 For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
-From Figure 1, we can clearly see that Algorithm 1 is significantly faster for highly-sparse instances, strongly outperforming even a state-of-the-art accelerated method, while as ${\mathbf{x}}^{*}$ becomes less-sparse, this advantage vanishes.
-(a) $\textrm{nnz}({\mathbf{x}}^{*})=10$
-(b) $\textrm{nnz}({\mathbf{x}}^{*})=30$
-(c) $\textrm{nnz}({\mathbf{x}}^{*})=50$
-Figure 1: Comparison of Algorithm 1 and the V-FISTA algorithm for Problem (1).
+For the accelerated non-Euclidean method we set the length of each epoch (the parameter $K$ in Algorithm 3) to $K=\left\lceil{25\sqrt{\frac{\beta_{1}{}s}{\alpha_{2}}}}\right\rceil$ which gave the best results.
+As for our Algorithm 1, we implemented three variants, all of them use exact line-search to determine the parameter $\gamma_{t}$, but with different choice of step-size $\eta$: I.
+using the theoretical step-size $\eta$ in Theorem 1 which, as discussed in Section 3.1 amounts to setting $\eta=\frac{\alpha_{2}}{48\beta_{1}s}$, II.
+using a larger manually-tuned heuristic step-size $\eta=\frac{\alpha_{2}}{2\beta_{1}s}$, and III.
+using an automatically-tuned step-size, as discussed in Remark 1, by trying on each iteration all step-sizes $\eta=\frac{2^{i}\alpha_{2}}{48\beta_{1}s}$ for $i\in\{0,1,\dots,5\}$ and taking the one which results in the largest decrease in function value.
+Note that for $i=0$ we get the theoretical step-size and thus this method is guaranteed to converge with at least the rate established in Theorem 1.
+For all methods that require the knowledge of $s$ (an upper-bound on $\mathrm{nnz}({\mathbf{x}}^{*})$), which includes our algorithm and the restarted non-Euclidean accelerated gradient method, we set for simplicity $s=\mathrm{nnz}({\mathbf{x}}^{*})$.
+Let us comment on the results displayed in Figure 1.
+First, we see that our Algorithm 1, when implemented exactly according to theory and with a fixed step-size, does not seem to work well in practice (perhaps with the exception of the case $\mathrm{nnz}({\mathbf{x}}^{*})=10$).
+This seems to be due to the small theoretical step-size (note the $1/48$ factor in the theoretical choice of $\eta$), which may be an artifact of our analysis.
+Once we use a manually-tuned step-size or the automatically-tuned step-size (which as stated above has the theoretical guarantee of Theorem 1), we observe state-of-the-art performance, clearly outperforming all benchmarks.
+Somewhat surprisingly, the restarted non-Euclidean accelerated gradient method fails to substantially outperform the Euclidean V-FISTA method, despite having a theoretically superior convergence rate when $\mathbf{x}$$^{*}$ is sparse.
+Similarly, except for the highly-sparse setting $\mathrm{nnz}({\mathbf{x}}^{*})=10$, the away-steps Frank-Wolfe also fails to outperform V-FISTA.
+(a) $\mathrm{nnz}({\mathbf{x}}^{*})=10$
+(b) $\mathrm{nnz}({\mathbf{x}}^{*})=40$
+(c) $\mathrm{nnz}({\mathbf{x}}^{*})=100$
+Figure 1: Empirical convergence of various algorithms for Problem (1).
 5 Open Questions
 We conclude with two open questions which we believe are of interest.
-First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems.
+First, it is interesting if our approach could be extended to also handle norm-regularized problems instead of norm-constrained problems.
 Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
@@
 We would like to thank Amir Beck and Marc Teboulle for fruitful discussions during the early stages of this work.
-Appendix A Proof of Lemma 1
+Appendix A Away-Steps Frank-Wolfe for Problem (1)
+In this section we show that the away-steps Frank-Wolfe method [14, 20] has a linear convergence rate of the form $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$ for Problem (1), assuming all optimal solutions lie on the boundary of the $\ell_{1}$ ball.
+Theorem 2 (Away-Steps Frank-Wolfe for Problem (1)).
+Consider Problem (1) and suppose that for any optimal solution $\mathbf{x}$$^{*}$ it holds that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$.
+Then, the Away-steps Frank-Wolfe algorithm (Algorithm 1 in [20], see also [14]) finds an $\epsilon$-approximated solution (in function value) after at most $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$ iterations.
+In particular, when $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$ (${\mathbf{A}}\succeq 0$), the overall runtime is $O\left({\frac{ns\beta_{1}R^{2}}{\alpha_{2}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$.
 Proof.
-Denote $k=\textrm{nnz}({\mathbf{x}})$.
+The theorem follows directly by instantiating Corollary 3.33 in [5], where we take the underlying norm to be the $\ell_{1}$ norm, and by using the fact that, as noted in [5], the $\ell_{1}$-pyramidal width of the $\ell_{1}$ ball is $1$.
+This leads to an $O\left({\frac{\beta_{1}R^{2}}{\alpha_{1}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$ upper-bound on the number of iterations to reach $\epsilon$-error, where $\alpha_{1}$ is the quadratic growth constant w.r.t.
+the $\ell_{1}$-norm.
+The proof follows since the assumption that all optimal solutions are on the boundary of the ball together with Lemma 1, implies that $\alpha_{1}\geq\frac{\alpha_{2}}{4s}$.
+∎
+Appendix B Accelerated Rates with Dense Updates — Recovering the rates in [15]
+In this section we give a concise description of the accelerated linear convergence rates reported in [15], which scale only with $\sqrt{\frac{\beta_{1}{}s}{\alpha_{2}}}$, but which rely on non-sparse updates.
+While the methods in [15] were developed for stochastic optimization, here we are concerned with a purely deterministic setting.
+For ease of presentation, throughout this section we will focus on Problem (1) (optimization over $\ell_{1}$ balls in $\mathbb{R}^{n}$) only, however the extension to Problem (2) (optimization over a matrix nuclear norm ball) is straightforward by replacing the entrywise $\ell_{1}$ norm with the matrix nuclear norm.
+The method is a straightforward combination of a restarted-accelerated non-Euclidean proximal gradient method due to Nesterov [23] with sparse projection steps of the iterates.
+Throughout this section we let $\mathcal{B}_{1}(R)$ denote the origin-centered $\ell_{1}$-ball of radius $R$ in $\mathbb{R}^{n}$.
+For ease of presentation, here (and only here) we assume that Problem (1) admits a unique optimal solution $\mathbf{x}$$^{*}$ (this is also assumed in [15]) and that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, however, here we do not require that the optimal solutions lie on the boundary of the $\ell_{1}$ ball.
+The following result is well-known, see for instance [15] .
+Lemma 3.
+Given ${\mathbf{x}}_{0}\in\mathcal{B}_{1}(R)$, consider the prox function
+$d({\mathbf{x}}):=\frac{1}{2}e\log{n}\cdot{}n^{(p-1)(2-p)/p}\|{{\mathbf{x}}-{\mathbf{x}}_{0}}\|_{p}^{2}\quad\textrm{for}\quad p=1+1/\log{n}.$ (9)
+Then, $d({\mathbf{x}})$ is $1$-strongly convex over $\mathcal{B}_{1}(R)$ w.r.t.
+the norm $\|{\cdot}\|_{1}$, and for any ${\mathbf{y}}\in\mathcal{B}_{1}(R)$, $d({\mathbf{y}})\leq\frac{e^{2}\log{n}}{2}\|{{\mathbf{y}}-{\mathbf{x}}_{0}}\|_{1}^{2}$.
+Algorithm 2 Nesterov’s accelerated gradient method (see Method 3.11 in [23])
+1: input: initialization point ${\mathbf{x}}_{0}\in\mathcal{B}_{1}(R)$, $\ell_{1}$-smoothness parameter $\beta_{1}$
+2: define the function $d({\mathbf{z}})$ as in Eq. (9) and set $\sigma=1$
+3:for $k=0,1\dots$ do
+4:  ${\mathbf{y}}_{k}\leftarrow\arg\min_{{\mathbf{y}}\in\mathcal{B}_{1}(R)}\langle{{\mathbf{y}}-{\mathbf{x}}_{k},\nabla{}f({\mathbf{x}}_{k})}\rangle+\frac{\beta_{1}}{2}\|{{\mathbf{y}}-{\mathbf{x}}_{k}}\|_{1}^{2}$
+5:  ${\mathbf{z}}_{k}\leftarrow\arg\min_{{\mathbf{z}}\in\mathcal{B}_{1}(R)}\left\{{\frac{\beta_{1}}{\sigma}d({\mathbf{z}})+\sum_{i=0}^{k}\frac{i+1}{2}\langle{{\mathbf{z}}-{\mathbf{x}}_{i},\nabla{}f({\mathbf{x}}_{i})}\rangle}\right\}$
+6:  ${\mathbf{x}}_{k+1}\leftarrow\frac{2}{k+3}{\mathbf{z}}_{k}+\frac{k+1}{k+3}{\mathbf{y}}_{k}$
+7:end for
+Theorem 3 (Theorem 2 in [23]).
+When applied to Problem (1), the sequence $({\mathbf{y}}_{k})_{k\geq 0}$ produced by Algorithm 2 satisfies
+$\forall k\geq 0:\qquad f({\mathbf{y}}_{k})-f({\mathbf{x}}^{*})\leq\frac{4\beta_{1}d({\mathbf{x}}^{*})}{(k+1)(k+2)}.$
+Algorithm 3 Restarted accelerated gradient method
+1: input: sparsity parameter $s\in[n]$, epoch length $K$
+[…]
+3:for $t=0,1\dots$ do
+4:  ${\mathbf{x}}_{t+1}\leftarrow$ the point ${\mathbf{y}}_{K}$ computed by Algorithm 2 after $K$ iterations, when initialized with the point $\widehat{{\mathbf{x}}}_{t}=\arg\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\mathrm{nnz}({\mathbf{x}})\leq s}\|{{\mathbf{x}}-{\mathbf{x}}_{t}}\|_{2}$ {i.e., $\widehat{{\mathbf{x}}}_{t}$ equals ${\mathbf{x}}_{t}$ on the $s$ largest (in absolute value) entries, and zero elsewhere}
+5:end for
+Theorem 4.
+Consider applying Algorithm 3 to Problem (1) with $K\geq\sqrt{\frac{64s\beta_{1}e^{2}\ln(n)}{\alpha_{2}}}-1$. Then,
+$\forall t\geq 0:\qquad f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})\leq\frac{\alpha_{2}R^{2}}{s}2^{-(t+3)}.$
+Thus, to reach an $\epsilon$-approximated solution (in function value) the number of calls to the first-order oracle of $f$ is $O\left({\sqrt{\frac{\beta_{1}{}s\ln(n)}{\alpha_{2}}}\ln\left({\frac{\alpha_{2}R^{2}}{s\epsilon}}\right)}\right)$.
+Proof.
+Fix some iteration $t\geq 0$ of Algorithm 3.
+Let $d_{t}({\mathbf{x}})$ denote the function $d({\mathbf{x}})$, as defined in Eq.
+(9), with ${\mathbf{x}}_{0}=\widehat{{\mathbf{x}}}_{t}$ — the initialization point in iteration $t$ of Algorithm 3.
+$\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t+1}}\|_{1}^{2}$ $\underset{(a)}{\leq}2s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t+1}}\|_{2}^{2}$
+$\leq 4s\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t+1}}\|_{2}^{2}+4s\|{{\mathbf{x}}_{t+1}-\widehat{{\mathbf{x}}}_{t+1}}\|_{2}^{2}$
+$\underset{(b)}{\leq}8s\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t+1}}\|_{2}^{2}\underset{(c)}{\leq}\frac{16s}{\alpha_{2}}\left({f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})}\right)$
+$\underset{(d)}{\leq}\frac{64s\beta_{1}d_{t}({\mathbf{x}}^{*})}{\alpha_{2}(K+1)^{2}}\underset{(e)}{\leq}\frac{32s\beta_{1}e^{2}\ln(n)\cdot{}\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}}{\alpha_{2}(K+1)^{2}},$
+where (a) holds since $\mathbf{x}$$^{*}$ is $s$-sparse by assumption and $\widehat{{\mathbf{x}}}_{t+1}$ is $s$-sparse by construction, (b) holds by definition of $\widehat{{\mathbf{x}}}_{t+1}$, (c) holds due to the $\ell_{2}$-quadratic growth assumption, (d) follows from Theorem 3 (when applied with the initialization point $\widehat{{\mathbf{x}}}_{t}$ and as a consequence the prox function is $d_{t}(\cdot)$ as defined above), and (e) follows from upper-bounding $d_{t}({\mathbf{x}}^{*})$ using Lemma 3.
+Thus, for $K\geq\sqrt{\frac{64s\beta_{1}e^{2}\log(n)}{\alpha_{2}}}-1$, and noting that $\widehat{{\mathbf{x}}}_{0}={\mathbf{x}}_{0}$, we can deduce that
+$\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}\leq 2^{-t}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{0}}\|_{1}^{2}.$
+Using Theorem 3 and Lemma 3 again, this yields
+$f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})$ $\leq\frac{4\beta_{1}d_{t}({\mathbf{x}}^{*})}{(K+1)^{2}}\leq\frac{2\beta_{1}e^{2}\log(n)\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}}{(K+1)^{2}}$
+$\leq\frac{\alpha_{2}}{s}2^{-(t+5)}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{0}}\|_{1}^{2}\leq\frac{\alpha_{2}R^{2}}{s}2^{-(t+3)}.$
+Remark 4.
+In the derivations above, for the sake of being concrete, we chose to present Nesterov’s Method 3.11 from [23] as Algorithm 2.
+A different choice could have been Method 5.6 from the same paper for which Theorem 3 still holds, however, the practical implementation of Method 5.6 might be easier since it requires that both prox computations are w.r.t.
+the same function $d(\cdot)$, instead of one w.r.t.
+the $\ell_{1}$ norm and one w.r.t.
+the function $d(\cdot)$ as in Algorithm 2.
+Appendix C Proof of Lemma 1
+Proof.
+Denote $k=\mathrm{nnz}({\mathbf{x}})$.
 Let $S\subseteq[n]$ be the set of indices for non-zero entries in $\mathbf{x}$}.
@@
 $\leq 2\sqrt{k}\|{{\mathbf{x}}-{\mathbf{y}}}\|_{2}.$
-Appendix B Proof of Lemma 2
+Appendix D Proof of Lemma 2
 Proof.
-Denote $k=\textrm{rank}({\mathbf{X}})$.
-Let ${\mathbf{P}}\in\mathbb{S}^{n}$ denote the projection matrix onto the subspace $\textrm{Im}({\mathbf{X}})$ and let ${\mathbf{P}}_{\perp}\in\mathbb{S}^{n}$ denote the projection matrix onto the orthogonal complement of $\textrm{Im}({\mathbf{X}})$.
-Throughout the proof we let $\|{\cdot}\|_{1}$ and $\|{\cdot}\|_{2}$ denote the trace norm (sum of singular values) and the Euclidean (Frobenius) norm in some space of real (not necessarily symmetric) matrices, respectively.
-We begin by recalling two important facts.
-First, for any real (not necessarily symmetric) matrix ${\mathbf{A}}$ it holds that
-$\|{{\mathbf{A}}}\|_{1}\leq\sqrt{\textrm{rank}({\mathbf{A}})}\|{{\mathbf{A}}}\|_{2}.$ (9)
-Second, for any two real (not necessarily symmetric) matrices of appropriate dimension (such that the multiplication is defined) ${\mathbf{A}},{\mathbf{B}}$ it holds that
-$\|{{\mathbf{A}}{\mathbf{B}}}\|_{2}\leq\|{{\mathbf{A}}}\|_{\infty}\|{{\mathbf{B}}}\|_{2},$ (10)
-where we let $\|{{\mathbf{A}}}\|_{\infty}$ denote the largest singular value of ${\mathbf{A}}$.
-Let us denote ${\mathbf{D}}={\mathbf{X}}-{\mathbf{Y}}$. It holds that,
-$\|{{\mathbf{D}}}\|_{1}$ $=\|{({\mathbf{P}}+{\mathbf{P}}_{\perp}){\mathbf{D}}({\mathbf{P}}+{\mathbf{P}}_{\perp})}\|_{1}$
-$\leq\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}+\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}+2\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}.$ (11)
-We continue by upper-bounding each of the three terms in the RHS separately.
-Note that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}})\leq k$. Thus, using Eq. (9) and then Eq. (10), and noting that $\|{{\mathbf{P}}}\|_{\infty}=1$, we have that
-$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (12)
-To bound the second term in the RHS of (B) we have,
-$\|{{\mathbf{P}}_{\perp}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}$ $=\|{{\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}}\|_{1}\underset{(a)}{=}\textrm{Tr}({\mathbf{P}}_{\perp}{\mathbf{Y}})=\textrm{Tr}(({\mathbf{I}}-{\mathbf{P}}){\mathbf{Y}})$
-$\underset{(b)}{\leq}\textrm{Tr}({\mathbf{X}})-\textrm{Tr}({\mathbf{P}}{\mathbf{Y}})=\textrm{Tr}({\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}})\leq\|{{\mathbf{P}}({\mathbf{X}}-{\mathbf{Y}}){\mathbf{P}}}\|_{1}$
-$\underset{(c)}{\leq}\sqrt{k}\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{2},$ (13)
-where (a) holds since ${\mathbf{Y}}$ is positive semidefinite, and thus so is ${\mathbf{P}}_{\perp}{\mathbf{Y}}{\mathbf{P}}_{\perp}$, (b) holds by the assumption that $\textrm{Tr}({\mathbf{Y}})=\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}=\textrm{Tr}({\mathbf{X}})$, and (c) is due to Eq.
-(12)
-Finally, for the last term in the RHS of (B), noting that $\textrm{rank}({\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp})\leq k$, using Eq.
-(9) and then Eq.
-(10) (and noting again that $\|{{\mathbf{P}}}\|_{\infty}=\|{{\mathbf{P}}_{\perp}}\|_{\infty}=1$), we have that
-$\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{1}\leq\sqrt{k}\|{{\mathbf{P}}{\mathbf{D}}{\mathbf{P}}_{\perp}}\|_{2}\leq\sqrt{k}\|{{\mathbf{D}}}\|_{2}.$ (14)
-The lemma follows from plugging-in (12), (B), and (14) into the RHS of (B). ∎
-Appendix C Sparse Projections onto the Bounded Positive Semidefinite Cone
+Throughout the proof we let $\|{\cdot}\|_{1}$, $\|{\cdot}\|_{2}$, $\|{\cdot}\|_{\mathrm{op}}$ denote the nuclear norm, the Euclidean (Frobenius) norm, and operator norm (largest singular value) in $\mathbb{R}^{m\times n}$, respectively.
+Denote ${\mathbf{D}}:={\mathbf{Y}}-{\mathbf{X}}$ and $r=\mathrm{rank}({\mathbf{X}})$.
+Let ${\mathbf{X}}={\mathbf{U}}{\mathbf{S}}{\mathbf{V}}^{\top}$ be a compact SVD of $\mathbf{X}$}.
+Define the orthogonal projectors ${\mathbf{P}}_{{\mathbf{U}}}:={\mathbf{U}}{\mathbf{U}}^{\top},{\mathbf{P}}_{{\mathbf{V}}}:={\mathbf{V}}{\mathbf{V}}^{\top}$, and accordingly define the linear maps $\mathcal{P}_{T},\mathcal{P}_{\perp}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}$ as
+$\mathcal{P}_{T}({\mathbf{Z}}):={\mathbf{P}}_{{\mathbf{U}}}{\mathbf{Z}}+{\mathbf{Z}}{\mathbf{P}}_{{\mathbf{V}}}-{\mathbf{P}}_{{\mathbf{U}}}{\mathbf{Z}}{\mathbf{P}}_{{\mathbf{V}}},\qquad\mathcal{P}_{\perp}({\mathbf{Z}}):=({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{U}}})\,{\mathbf{Z}}\,({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{V}}}).$
+Write
+${\mathbf{D}}_{T}:=\mathcal{P}_{T}({\mathbf{D}}),\qquad{\mathbf{D}}_{\perp}:=\mathcal{P}_{T^{\perp}}({\mathbf{D}}),$
+so that ${\mathbf{D}}={\mathbf{D}}_{T}+{\mathbf{D}}_{\perp}$.
+Clearly, we have that
+$\|{{\mathbf{D}}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}+\|{{\mathbf{D}}_{\perp}}\|_{1}.$ (10)
+Additionally, since ${\mathbf{U}},{\mathbf{V}}$ are of rank $r$, it follows from the definition of the map $\mathcal{P}_{T}$ that $\mathrm{rank}({\mathbf{D}}_{T})\leq 2r$. Thus, we have that
+$\|{{\mathbf{D}}_{T}}\|_{1}\leq\sqrt{\mathrm{rank}({\mathbf{D}}_{T})}\|{{\mathbf{D}}_{T}}\|_{2}\leq\sqrt{2r}\|{{\mathbf{D}}_{T}}\|_{2}.$ (11)
+Thus, given (10) and (11), to prove the lemma, it suffices to show that I.
+$\|{{\mathbf{D}}_{\perp}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}$ and II.
+that $\|{{\mathbf{D}}_{T}}\|_{2}\leq\|{{\mathbf{D}}}\|_{2}$.
+To see why $\|{{\mathbf{D}}_{T}}\|_{2}\leq\|{{\mathbf{D}}}\|_{2}$ holds, it suffices to show that $\mathcal{P}_{T}$ is an orthogonal projection, meaning it is a contraction in Frobenius norm.
+To see this, observe that by definition, for any $\mathbf{Z}$}, $\mathcal{P}_{\perp}(\mathcal{P}_{\perp}({\mathbf{Z}}))=\mathcal{P}_{\perp}({\mathbf{Z}})$, meaning $\mathcal{P}_{\perp}$ is an orthogonal projection.
+Since, by definition $\mathcal{P}_{T}=\mathcal{I}-\mathcal{P}_{\perp}$, where $\mathcal{I}$ is the identity map, it follows that $\mathcal{P}_{T}$ is also an orthogonal projection.
+Thus, in the remainder of the proof we show that $\|{{\mathbf{D}}_{\perp}}\|_{1}\leq\|{{\mathbf{D}}_{T}}\|_{1}$.
+Write the compact SVD of ${\mathbf{D}}_{\perp}=\widetilde{\mathbf{U}}\,\widetilde{\mathbf{S}}\,\widetilde{\mathbf{V}}^{\top}$.
+Since ${\mathbf{D}}_{\perp}=({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{U}}}){\mathbf{D}}({\mathbf{I}}-{\mathbf{P}}_{{\mathbf{V}}})$, we have that
+${\mathbf{U}}^{\top}\widetilde{\mathbf{U}}=\mathbf{0},\qquad{\mathbf{V}}^{\top}\widetilde{\mathbf{V}}=\mathbf{0}.$ (12)
+Define ${\mathbf{W}}:=\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}$.
+Note that $\|{\mathbf{W}}\|_{\mathrm{op}}=1$.
+Note that due to (12) we have that $\|{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}}}\|_{\mathrm{op}}=1$.
+Using the variation characterization of the nuclear norm, we have that
+$\|{{\mathbf{Y}}}\|_{1}$ $=\|{{\mathbf{X}}+{\mathbf{D}}}\|_{1}$
+$=\max_{\|{{\mathbf{Q}}}\|_{\mathrm{op}}\leq 1}\langle{{\mathbf{Q}},{\mathbf{X}}+{\mathbf{D}}}\rangle$
+$\underset{(a)}{\geq}\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{X}}+{\mathbf{D}}}\rangle$
+$=\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{X}}}\rangle+\langle{{\mathbf{W}},{\mathbf{X}}}\rangle+\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{D}}}\rangle$
+$\underset{(b)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}},{\mathbf{D}}_{T}+{\mathbf{D}}_{\perp}}\rangle$
+$\underset{(c)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle+\langle{{\mathbf{W}},{\mathbf{D}}_{\perp}}\rangle$
+$\underset{(d)}{=}\|{{\mathbf{X}}}\|_{1}+\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle+\|{{\mathbf{D}}_{\perp}}\|_{1},$
+where (a) holds since, as noted above, $\|{{\mathbf{U}}{\mathbf{V}}^{\top}+{\mathbf{W}}}\|_{\mathrm{op}}=1$, (b) holds since using the SVD of $\mathbf{X}$} we have that $\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{X}}}\rangle=\mathrm{Tr}({\mathbf{U}}{\mathbf{V}}^{\top}{\mathbf{X}}^{\top})=\mathrm{Tr}({\mathbf{U}}{\mathbf{V}}^{\top}{\mathbf{V}}{\mathbf{S}}{\mathbf{U}}^{\top})=\mathrm{Tr}({\mathbf{S}})=\|{{\mathbf{X}}}\|_{1}$, (c) holds since by definition of $\mathbf{D}$$_{\perp}$ we have that $\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{\perp}}\rangle=0$ and by definition of ${\mathbf{D}}_{T}$ and $\mathbf{W}$} we have that $\langle{{\mathbf{W}},{\mathbf{D}}_{T}}\rangle=0$, and (d) holds since, by definition of $\mathbf{W}$} and using the SVD of $\mathbf{D}$$_{\perp}$, we have that $\langle{{\mathbf{W}},{\mathbf{D}}_{\perp}}\rangle=\mathrm{Tr}(\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}{\mathbf{D}}_{\perp}^{\top})=\mathrm{Tr}(\widetilde{\mathbf{U}}\widetilde{\mathbf{V}}^{\top}\widetilde{\mathbf{V}}\widetilde{\mathbf{S}}\widetilde{\mathbf{U}}^{\top})=\mathrm{Tr}(\widetilde{\mathbf{S}})=\|{{\mathbf{D}}_{\perp}}\|_{1}$.
+Using the assumption that $\|{{\mathbf{Y}}}\|_{1}\leq\|{{\mathbf{X}}}\|_{1}$ and the fact that $\|{{\mathbf{U}}{\mathbf{V}}^{\top}}\|_{\mathrm{op}}=1$, we indeed have that
+$\|{{\mathbf{D}}_{\perp}}\|_{1}\leq-\langle{{\mathbf{U}}{\mathbf{V}}^{\top},{\mathbf{D}}_{T}}\rangle\leq\|{{\mathbf{U}}{\mathbf{V}}^{\top}}\|_{\mathrm{op}}\cdot\|{{\mathbf{D}}_{T}}\|_{1}=\|{{\mathbf{D}}_{T}}\|_{1},$
+which concludes the proof. ∎
+Appendix E Sparse Projections onto the Nuclear Norm Ball
 Lemma 4.
-Let ${\mathbf{Z}}\in\mathbb{S}^{n}$ with eigen-decomposition ${\mathbf{Z}}=\sum_{i=1}^{n}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ ($\lambda_{1}\geq\lambda_{2}\geq...$) and $s\in[n]$.
-A Euclidean projection of ${\mathbf{X}}$ onto $\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,~\textrm{Tr}({\mathbf{X}})\leq R,\textrm{rank}({\mathbf{X}})\leq s\}$ is given by the matrix ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\lambda_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$, where $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of the vector $(\lambda_{1},\dots,\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
+Let ${\mathbf{Z}}\in\mathbb{R}^{m\times n}$ with SVD ${\mathbf{Z}}=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}{\mathbf{u}}_{i}{\mathbf{v}}_{i}^{\top}$ ($\sigma_{1}\geq\sigma_{2}\geq$...) and $s\in[\min\{m,n\}]$.
+A Euclidean projection of $\mathbf{Z}$} onto $\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,\mathrm{rank}({\mathbf{X}})\leq s\}$ is given by ${\mathbf{Z}}^{\prime}=\sum_{i=1}^{s}\sigma_{i}^{\prime}{\mathbf{u}}_{i}{\mathbf{v}}_{i}^{\top}$, where $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$ is the Euclidean projection of the vector $(\sigma_{1},\dots,\sigma_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
 Proof.
-[…]
+Given some ${\mathbf{Z}}\in\mathbb{R}^{m\times n}$ and $s\in[\min\{m,n\}]$, let $\mathbf{Z}$$^{\prime}$ be as described in the lemma, and let ${\mathbf{Y}}\in\{{\mathbf{X}}\in\mathbb{R}^{m\times n}~|~\|{{\mathbf{X}}}\|_{1}\leq R,~\mathrm{rank}({\mathbf{X}})\leq s\}$ given by its SVD ${\mathbf{Y}}=\sum_{i=1}^{s}\gamma_{i}{\mathbf{p}}_{i}{\mathbf{w}}_{i}^{\top}$, which clearly implies that $(\gamma_{1},\dots,\gamma_{s})\in\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
+Recall $\|{\cdot}\|$ denotes the Euclidean norm.
+It holds that
 $\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}$ $=\|{{\mathbf{Y}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}}\|_{2}^{2}-2\langle{{\mathbf{Y}}-{\mathbf{Z}}^{\prime},{\mathbf{Z}}}\rangle$
-$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i}$
-$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\lambda_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\lambda_{i}+2\sum_{i=1}^{s}\lambda_{i}^{\prime}\lambda_{i},$
-where the last inequality is due to von Neumann’s trace inequality.
+$=\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\sigma_{j}^{\prime 2}-2\langle{{\mathbf{Y}},{\mathbf{Z}}}\rangle+2\sum_{i=1}^{s}\sigma_{i}^{\prime}\sigma_{i}$
+$\geq\sum_{i=1}^{s}\gamma_{i}^{2}-\sum_{j=1}^{s}\sigma_{j}^{\prime 2}-2\sum_{i=1}^{s}\gamma_{i}\sigma_{i}+2\sum_{i=1}^{s}\sigma_{i}^{\prime}\sigma_{i},$
+where the last inequality is due to von Neumann’s trace inequality [21].
 Rearranging, we have that
-$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\lambda_{i})^{2}-\sum_{j=1}^{s}(\lambda_{j}^{\prime}-\lambda_{j})^{2}\geq 0,$
-where the last inequality follows since $(\lambda_{1}^{\prime},\dots,\lambda_{s}^{\prime})$ is the Euclidean projection of $(\lambda_{1},\dots\lambda_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
+$\|{{\mathbf{Y}}-{\mathbf{Z}}}\|_{2}^{2}-\|{{\mathbf{Z}}^{\prime}-{\mathbf{Z}}}\|_{2}^{2}\geq\sum_{i=1}^{s}(\gamma_{i}-\sigma_{i})^{2}-\sum_{j=1}^{s}(\sigma^{\prime}_{j}-\sigma_{j})^{2}\geq 0,$
+where the last inequality holds since $(\sigma_{1}^{\prime},\dots,\sigma_{s}^{\prime})$ is the Euclidean projection of $(\sigma_{1},\dots\sigma_{s})$ onto $\{{\mathbf{x}}\in\mathbb{R}^{s}~|~{\mathbf{x}}\geq 0,~\sum_{i=1}^{s}{\mathbf{x}}_{i}\leq R\}$.
 ∎